### v1

# Universal Vessel Segmentation for Multi-Modality Retinal Images

Abstract
We identify two major limitations in the existing studies on retinal vessel segmentation: (1) Most existing works are restricted to one modality, i.e, the Color Fundus (CF). However, multi-modality retinal images are used every day in the study of retina and retinal diseases, and the study of vessel segmentation on the other modalities is scarce; (2) Even though a small amount of works extended their experiments to limited new modalities such as the Multi-Color Scanning Laser Ophthalmoscopy (MC), these works still require finetuning a separate model for the new modality. And the finetuning will require extra training data, which is difficult to acquire. In this work, we present a foundational universal vessel segmentation model (UVSM) for multi-modality retinal images. Not only do we perform the study on a much wider range of modalities, but also we propose a universal model to segment the vessels in all these commonly-used modalities. Despite being much more versatile comparing with existing methods, our universal model still demonstrates comparable performance with the state-of-the-art finetuned methods. To the best of our knowledge, this is the first work that achieves cross-modality retinal vessel segmentation and also the first work to study retinal vessel segmentation in some novel modalities.  [^1] Our codes and model will be available at the time of publication. We are also actively looking for approaches to release the three new vessel segmentation datasets. However, since retinal images and retinal vessels contain sensitive personal information, we can not make promise at this moment.
Index Terms:
retina, retinal images, vessel segmentation, image translation, topology-preserving image segmentation.
I Background and Introduction
Retinal vessel segmentation is a fundamental task in retinal image processing. It has multiple useful applications, including retinal image registration [1, 2], human biometric identification [3], retinal artery-vein classification [4] and tree topology analysis in digital images [5]. Clinically, the evaluation of retinal vessels is also important since the vessel features are predictive in several systemic diseases [6, 7, 8] and vessel segmentation is the key towards the automation of the evaluation. However, most works in retinal vessel segmentation focus on only the Color Fundus (CF) images. CF uses white visible light to illuminate the retina and the reflected light is used for generating the image. Nevertheless, although CF was the most widely-used modality in the past, it requires dilation before the imaging and is uncomfortable for the patients. Additionally, the pathological information provided by CF is limited. Therefore, in more recent years, CF is less and less used and emerging new imaging techniques become increasingly popular. These modalities include (Fig. 1): (1) Multi-Color Scanning Laser Ophthalmoscopy (MC), which uses more intense laser beam and allow exposing structures in deeper layers of the retina; (2) Florescence Angiography (FA), which uses florescence dye to illuminate the retinal vessels and are commonly used in the diagnosis of diseases related to neovascularization; (3) Fundus AutoFlorescence (FAF), which does not rely on external light source but use natural fluorescence of the retina and provide insights into the health and metabolism of the retinal pigment epithelium (RPE) and photoreceptor cells; (4) Infrared/Near-Infrared Reflectance (IR/NIR), which uses Infrared or Near-Infrared lights to illuminate the image and allows seeing deeper features on the retina such as RPE abnormalities. (5) Optical Coherence Tomograph Angiograph (OCTA), which uses low-coherence interferometry to measure the backscattered light from retinal layers. It is primarily used to image the micro-vascular structure on the retina. Nevertheless, OCTA has a much smaller imaging area than the other modalities and has significantly different retinal vessel morphology and topology. Therefore, in this work we focus on the first four modalities and the CF.
Figure 1: Overview of multi-modality retinal images and the results from our UVSM. In each image, the upper half is the original image and the lower half is the segmentation. In the middle, the camera (left) and modality (right) of the image is labeled. ‘Heidel’ refers to Heidelberg Spectralis.
One major challenge that impedes the development of machine learning-based retinal vessel segmentation algorithms is the lack of data. Due to the complexity of the retinal vessels and the existence of many fine details, labeling is very time-consuming. Existing retinal vessel segmentation datasets (especially considering high label quality) are scarce and small even for the CF modality, which usually only have 20-40 images. In addition, to the best of our knowledge, there are no publicly available vessel segmentation datasets for FAF and IR. For FA, there is only a very small public dataset with 8 images [9]. In this work, we prepare three high-quality datasets for FA, FAF and IR, respectively. Each dataset has 40 images and the vessels are professionally and carefully labeled. Nevertheless, despite the availability of the new datasets, in this work, we only use them for evaluation purpose. The reason is that: we believe that there can never be enough annotated images to allow training a powerful universal model. Instead, we need to train the model with limited existing data and seek to overcome the challenge through innovations in the algorithm. In fact, our segmentation network is trained with only 19 annotated CF images from the widely-used DRIVE [10] dataset.
There are limited previous works which are extended beyond CF modality [11, 12, 13]. However, not only they did not work with as comprehensive modalities as we do, but also in these work one still needs to finetune a separate model on the target modality and each model only works for one modality. We instead aim to build a universal model to use a single model and allows robust segmentation for all the commonly-used modalities. The model will serve as an important tool in cross-modality retinal image study, which becomes increasingly important in modern research of the retina. To achieve the universal segmentation with limited labeled data. We propose our model in two major steps (Fig. 2). Firstly, we propose to use image translation as a preliminary domain adaptation method and train a model that translate input images from arbitrary modalities and cameras into a uniform Topcon (The Topcon TRC-50DX series is explicitly referred to as the gold standard in color fundus photography) CF image modality. We accessed abundant of de-identified multi-modality retinal images from multiple widely-used cameras (Topcon, Heidelberg Spectralis, Optos) at the Jacobs Retina Center at Shiley Eye Institute, University of California, San Diego (UCSD). Those images allow us to train a useful image translation model in a self-supervised approach. Furthermore, we adopt data augmentation to account for the slight domain difference with the other CF cameras which our center does not have, including Canon and Zeiss. Secondly, we train a segmentation model on the Topcon CF domain to perform the segmentation. However, image translation can not fully eliminate the domain gaps between different modalities given such a difficult many-to-one translation task. To solve this problem, we propose to use a topology-aware segmentation method that learns topological features of the vessels in addition to their conventional pixel-level features. The topology-aware segmentation methods were initially proposed to improve the topological accuracy of the segmentation of tubular structures within a certain type of image. We extend these methods to our task and found that not only do they improve the topological accuracy in the segmentation (e.g., preserving vessel continuity, branching and loop features), but also they force the segmentation to be topologically correct so when images with slight domain gap are inputted, the model can still make the correct segmentation.
Our universal vessel segmentation model is evaluated on 7 diverse datasets (including 3 new datasets we prepared for this work) of the 5 most commonly-used retinal image modalities (CF,FA,FAF,MC,IR/NIR) taken from 5 most commonly-used cameras (Topcon, Canon, Zeiss, Heidelberg Spectralis, Optos). Our universal model is compared with state-of-the-art methods which are finetuned on each of the 7 datasets, respectively, and is shown to achieve comparable performance. Extensive ablation studies were also performed to investigate how different design choices in the proposed pipeline affects the performance of the universal vessel segmentation. The main contributions of this paper can be summerized as:
  • We propose a foundational universal vessel segmentation model (UVSM) for multi-modality retinal images which can perform robust vessel segmentation for all the commonly-used modalities and cameras using one single model without providing any modality information.
  • We investigate image-to-image translation between retinal images from different modalities. We explore different Generative Adversarial Nets (GAN)-based and Diffusion Model-based image translation methods and evaluate their performance on retinal images.
  • We propose to use a topology-aware image segmentation method to train our segmentation model to allow more robust segmentation across modalities. Furthermore, we discover an additional merit of the topology-aware segmentation loss functions in segmenting images with domain gaps in addition to their conventional applications.
  • We collect and annotate three high-quality vessel segmentation datasets for the FA, FAF and IR retinal images, respectively.
  • We provide a refined version of the widely-used DRIVE dataset with better topological accuracy in the vessel labels (to be discussed in section IV-A2).
II Related Work
II-A Retinal Vessel Segmentation
Retinal vessel segmentation classifies each pixel on the retinal images into foreground (vessel) and background (non-vessel). And the literature focuses on segmenting vessels in the CF images. Various machine learning-based or non-machine learning-based methods were proposed. A majority of the early methods are based on the matched filter [14, 15, 16], which convolves retinal images with predefined or learned kernels and then adaptively threshold the image to obtain the vessel segmentation map. Another type of methods use tracking algorithms to mathematically estimate the growth of the retinal vessel trees from seed points [17, 18]. The segmentation can then be derived by extending around the estimated vessel centerlines using the estimated vessel diameters. Some methods also adopt morphological image processing methods such as the top-hat operation to extract the vessel or as a post-processing step [19, 20]. Furthermore, numerous traditional machine learning-based methods were also applied in retinal vessel segmentation. Many classic models were used, such as the support vector machine [21], random forest [22] and Adaboost [23] for supervised methods, and Gaussian mixture model [24], Fuzzy C-Means [25] for unsupervised methods. In more recent years, deep learning-based methods dominate the literature. They majorly rely on the Convolutional Neural Networks (CNN) as a backbone model due to the existence of vast fine details and the whole-spreading characteristic of the vessels. An introductory work is the Deep Retinal Image Understanding (DRIU) [26], which adopts a VGG [27] network structure as encoder and assign specialized output layers to the embedded features to segment the vessels and the optic disc from the retina. Based on a widely-used U-Net [28] network structure and to seek further improvement for retinal vessel segmentation, [29, 30, 31, 32, 33] proposed to add various global reasoning modules at the bottom of the network. [32, 33, 34, 35, 31] proposed to add feature aggregation modules in the residual connection at each level of the network. [36, 37] proposed to use deformable convolution where the convolution kernel shapes are learned to adapt to the elongated morphology of the retinal vessels. With similar motivation, [38] proposed an orientation-selective convolution to learn selective receptive fields of the kernel to adapt to the shape of the vessels. [31] proposed to use improved feature fusion methods to better preserve the fine details of vessels in the segmentation. In addition to a single U-Net-like segmentation network, several works [39, 38, 40] proposed to use cascaded networks or a single looped network to iteratively refine the segmentation.
II-B Unpaired Image-to-Image Translation
Image-to-image translation aims to transfer a source image to a target image domain while preserving its structural information. However, it is difficult to prepare paired retinal images from the source and target domain for training. Therefore, in this work we look into the unpaired image-to-image translation methods where the projection between the two domains can be learned with unpaired images with no structural correspondences. A representative and most-commonly used method in this task is CycleGAN [41], which introduces a cycle consistency that projects the translated image back to the source domain to force retaining the structural information in the translated image. Following works further introduce geometric consistency [42], mutual information regularization [43] and contrastive unpaired translation [44] to improve the computational efficiency over the cycle consistency-based methods during the training phrase. [45, 46] proposed one-to-many transfer schemes which allow translating images to multiple domains. While the above works are based on GAN backbone and achieved good image translation performance, diffusion models [47, 48, 49] have recently gained more research attention in this topic for easier training and superior image quality. However, the methods [50, 51, 52] are computationally intensive and were only applied on small images. For example, [52] combines diffusion model with adversarial learning and formulate the training in a cycle-consistency framework, which requires training two diffusion models at the same time. Most recently, [53] proposed an Unpaired Neuron Schrodinger Bridge (UNSB) with a novel scheme to simulate the Schrodinger Bridge, which is a random process that interpolates between two image domains, via a diffusion process. It has higher computational efficiency and makes application to higher-resolution images (e.g., retinal images) possible.
Image style transfer [54, 55, 56, 57, 58] is another method which can be seen as a special type of unpaired image-to-image translation but with a perceptual inductive bias. The majority of the methods use a perceptual feature space to compute a style loss and inject the style of a reference image to the input content image, while using another self-comparison content loss to retain the content information of the input in the transferred image. We also explore methods in this topic since style transfer is used for non-CF retinal vessel segmentation in [12, 13].
II-C Topology-Aware Image Segmentation
Topology-aware image segmentation aims to preserve the topological accuracy in the segmentation of tubular and net-like structures. The literature mainly researches a topological loss function that extracts topological features from the ground truth and segmentation and minimizes the difference between them. One type of methods extract certain features from the segmentation which are proximity to the actual topological features of the foreground objects. A pioneer work [59] uses an ImageNet-pretrained VGG [60, 27] network after the segmentation network. The perceptual features extracted from different layers of the pretrained VGG network is shown to be approximations of the topological features. The topological loss is then computed between the ground truth and prediction perceptual features. Another topological segmentation loss function [61] was proposed exclusively for the the segmentation of glands. The loss is computed according to the iterations needed to derive the gland skeletons by the image erosion operation. [62] proposed to approximate the topological features of tubular structures by a relationship between the foregrounds in the ground truth and prediction and their soft-skeletons. Using such a relationship as a loss function is shown to strengthen the connectivity along the tubular structure thus providing better topological accuracy.
Another type of methods use strict topological features of the segmentation which are based on the computational topology theory and receieved more research attentions in recent years. The topological features are mainly based on the persistent homology theory [63, 64, 65, 66, 37, 67, 68, 69]. A revolutionary work that first introduce this concept in image segmenation is [63], which computes the persistent features from the ground truth and prediction, matching the features using a Wasserstein distance [70] and use the distance as the loss function. The following works share similar idea. [64] also uses the Wasserstein distance but apply the method on 3D point cloud data. [67] proposed to match the longest persistent barcodes in the prediction to the ground truth. And [37] proposed to use a Hausdorff distance [71] to matches the persistent diagrams to improve the sensitivity of the matching to the outliers. [69] and [68] both aim to compute more precise persistent feature matching, as in earlier methods the actual matchings results are highly inaccurate due to only considering the global topological features. [69] and [68] instead allow the loss to act more selectively on the topological features thus improve the efficiency of the topological loss function. [68] proposed to use an Induced Matching [72] method which constructs a common ambient space between the ground truth and prediction and maps the persistent features in between to find a better matching. However, the computational cost of this method is very high, especially for large images. And the retinal images usually have high resolutions. Alternatively, [69] proposed to leverage the spatial locations of the persistent features referencing their critical cell locations for more faithful persistent feature matching and is computationally much more efficient than [68].
Figure 2: Pipeline for our proposed universal vessel segmentation model, which includes the 2-stage training and the inference.
III Universal Retinal Vessel Segmentation Model
III-A Overview
We discuss the overall training and inference pipeline (Fig. 2) in this sub-section and provide more details about the major components in the pipeline in the following sub-sections. Given multi-modality retinal images, we first adopt a simple yet widely-used pre-processing method that takes only the green channel from the RGB image. All of the training and inference are then performed on the green-channel images.
Following the pre-processing, we train an domain adaptation model for multi-modality retinal images. Although several domain adaptation methods were proposed for retinal vessel segmentation, such as using the pseudo-label methods [39], adversarial learning [33] and asymmetrical maximum classifier discrepancy [73], they were proposed only to address the domain gap between different datasets (from different cameras) in CF images. Moreover, these methods still require unsupervised finetuning on the target domain images, which deviate from our goal of universal segmentation. Therefore, in this work we propose to use image translation as the domain adaptation method, which allows us to train a single model for all the modalities.
The image translation model aims to translate images from all the modalities to a uniform Topcon CF modality, and is trained between equivalent number of non-CF images (FA, FAF, MC, IR/NIR) and the Topcon CF images, i.e., during the training, we treat all the non-CF modalities as one type of image. More information of the training dataset is provided in section IV-A1. Furthermore, the translation network is trained with an identity loss (will be discussed in detail in section III-B) which allows us to use a single translation network to transfer all the images to the Topcon CF domain without the need to specify whether they are CF or non-CF. In addition, the translation model also translates CF images from other cameras to Topcon.
After the image translation model is trained, we use it to translate our segmentation training set (19 images from the DRIVE [10] training set with vessel labels) to the Topcon CF domain. The translated images are combined with the original images as an augmented training set to train the downstream segmentation network. Nevertheless, since we are training a difficult many-to-one (FA, FAF, MC, IR/NIR to CF) translation, the translation model cannot perfectly address the domain gap between the non-CF images and the CF images (Fig. 12). We therefore propose to further solve this problem by introducing a topology-aware segmentation method. In addition to a conventional pixel-wise loss function (BCELoss), the segmentation network is jointly trained with two topological loss functions.
At inference, we can disregard the modality of the image. Given a retinal image from an arbitrary modality, we simply need to (1) extract the green channel; (2) apply the translation network to the green-channel image; (3) apply the topology-aware segmentation network to the translated image to obtain the final segmentation.
III-B Image Translation Network
Multiple GAN-based or Diffusion model-based image translation methods are explored for our task. For Diffusion-based method, we experiment with StyleDiffusion [58] and Unpaired Neuron Schrodinger Bridge (UNSB) [53]. For GAN-based method, we experiment with CycleGAN [41]. We will show in section IV-D2 that CycleGAN, despite being a classic method, outperforms the other two most recent methods in most datasets. Therefore, we pick CycleGAN as the backbone of our image translation model. Consequently, the following sections will be discussing the implementation based on CycleGAN and we refer the readers to [58, 53] for details about the other two methods and to section IV-D2 for our implementation details.
Figure 3: Overview of our image translation network, which is based on CycleGAN [41]. $G$ refers to the GAN generators and $D$ for the discriminators. The gray dotted lines illustrate how loss functions are computed between different images.
To train the translation network, we treat all the non-CF images (FA, FAF, MC, IR/NIR) as image type A and the Topcon CF images as image type B. At each iteration, a pair of image A batch and image B batch is drawn from the training set. As illustrated by Fig. 3, generator A is trained to translate all the type A images (non-CF) to a fake type B (CF) and generator B is used to translate type B to a fake type A, which is a neutral non-CF domain. Both generator A and B are learned with adversarial learning [74, 75], where the generator and discriminator are alternatively updated. At each iteration, given the image batch $I_{A}$, the generators (A+B) are first updated with the objectives (discriminators frozen):
$\textit{L}_{G_{A}}=\textit{L}_{adv.-G_{A}}+\lambda_{c}\textit{L}_{cycle-A}+\lambda_{i}\textit{L}_{identity-A}$ (1)
where the adversarial loss defined as:
$\textit{L}_{adv.-G_{A}}=\left|D_{A}(G_{A}(I_{A}))-1\right|^{2}$ (2)
Furthermore, the translated fake CF images are again fed into the opposite generator to reconstruct the non-CF images. The reconstruction is supervised by a cycle consistency loss:
$\textit{L}_{cycle-A}=\left|I_{A}-G_{B}(G_{A}(I_{A}))\right|$ (3)
and is controlled by a weight $\lambda_{c}$. Finally, the real images are also passed through their opposite generator to generate the image itself. This step is supervised with an identity loss:
$\textit{L}_{identity-A}=\left|I_{A}-G_{B}(I_{A})\right|$ (4)
and is controlled by a weight $\lambda_{i}$. The identity loss was initially proposed to correct the tint in the translated image [41]. However, we find it especially useful in our task since it allows us to use the generator $G_{A}$ to map the CF images to itself. As a result, at inference we only need to apply one generator $G_{A}$ to all the input images regardless of their modalities. If the image is non-CF, the translation model will convert it to CF. If the image is already CF, the output image will remain a CF image. It allows our universal model to work without the need to provide any modality information about the input.
After the generator is updated, the discriminator is updated with the adversarial loss (generators frozen):
$\textit{L}_{adv.-D_{A}}=\frac{1}{2}\left|D_{A}(I_{B})-1\right|^{2}+\frac{1}{2}\left|D_{A}(G_{A}(I_{A}))\right|^{2}$ (5)
For the opposite image batch $I_{B}$, the generators (B+A) and discriminator (B) are updated in the same way, as illustrated in Fig. 3. To be more specific, we first update the generators for both $I_{A}$ and $I_{B}$, and then update the discriminators.
III-C Topology-aware Segmentation Network
The downstream segmentation network is trained with a Perceptual Topological Loss $\mathcal{L}_{tc}$ and a Persistent Topological Loss $\mathcal{L}_{ts}$, introduced below accordingly.
III-C1 Perceptual Topological Loss
We first propose to use a perceptual topological loss [59] where the topological information is approximated by the perceptual features encoded by an ImageNet [60] pretrained VGG-16 [27] network. As illustrated by Fig. 4, image embeddings $\mathcal{E}_{l}(\cdot)$ are computed from the predicted vessel probability map $I_{pred}$ and the ground truth $I_{GT}$, where $l$ denotes different layers in the VGG-16 network. In practice, we take the output of the $relu2\_2$, $relu3\_3$ and $relu4\_3$ layers. Then, the perceptual topological loss function $\mathcal{L}_{tc}$ is defined as:
$\mathcal{L}_{tc}=\sum_{l}^{N_{l}}\left|\mathcal{E}_{l}(I_{pred})-\mathcal{E}_{l}(I_{GT})\right|^{2}$ (6)
Figure 4: Illustration for the perceptual topological loss function.
III-C2 Persistent Topological Loss
In 2D digital images, the topology is described by their 0- and 1-homology features, i.e., the connected components and the loops. The persistent homology is a method to describe a process where homology features are created and destroyed in a sequence (called ‘filtration’) of cubical complex [76] (which encodes an image and the higher-dimensional relationships between its pixels to structured cells) and its sub-complexes determined by the values of cells. Although it deviates from the definition and how persistent homology is computed, a simpler way to interpret persistent homology in 2D digital images is by thresholding the image from its highest pixel value to its lowest pixel value and stop at every pixel value that appears in the image. Each thresholded image is a binary image where we can determine the connected components and the loops inside. As demonstrated by the example in Fig. 5, when the threshold value decreases, new pixels are added to the binary image, leading to creation of new homology features or destruction of old features. A homology feature that ‘lives’ through the decreasing thresholds is then called a ‘persistent feature’. We refer the readers to [77] for more details and strict definitions.
Figure 5: A toy example visualizing persistent homology in images: (a) a gray-scale image; (b) the filtration visualized by thresholded binary images. Under each image, the digits represent the filtration value (threshold value), the number of 0-homology features and the number of 1-homology features; (c) the 0-persistent diagram; (d) the 1-persistent diagram.
The persistent homology of an image can be described by a persistent diagram, where each point on the diagram represent one persistent feature, whose coordinate is (1--Creation Value, 1--Destruction Value).
We propose to use the Spatial-Aware Topological Loss Function (SATLoss) [69] as the persistent topological loss function. After computing persistent diagrams $\mathcal{D}(I_{pred})$ and $\mathcal{D}(I_{GT})$, the persistent features are matched using a spatially-weighted Wasserstein distance [70]:
$W_{q}(\mathcal{D}(I_{pred}),\mathcal{D}(I_{GT}))=$ (7)
$[\inf_{\eta:\mathcal{D}(I_{pred})\rightarrow\mathcal{D}(I_{GT})}\sum_{p\in\mathcal{D}(I_{pred})}s_{p}\left\|p-\eta(p))\right\|^{q}]^{\frac{1}{q}}$
where $s_{p}=\left\|c_{b}(p)-c_{b}(\eta(p))\right\|^{q}$ is a spatial weight that is computed between the creation cells of the persistent features. $p$ and $\eta(p)$ are a persistent feature from $I_{pred}$ and a candidate persistent feature from $I_{GT}$ to be matched with. $q$ is the norm and in practice we take $q=2$. After finding the optimal matching by solving the optimal transport problem, the persistent topological loss function is computed as (reformulating $p$ and $\eta(p)$ by their creation and destruction values):
$\mathcal{L}_{ts}=\sum_{p\in\mathcal{D}(I_{pred})}s^{*}_{p}([b(p)-b(\eta^{*}(p))]^{2}+[d(p)-d(\eta^{*}(p))]^{2})$ (8)
where $\eta^{*}(p)$ is the optimal matching of $p$ in $\mathcal{D}(I_{GT})$ and $s^{*}_{p}$ is the corresponding spatial weight. $b(\cdot)$ and $d(\cdot)$ are the creation and destruction values of the topological features. The values are taken from $I_{pred},I_{GT}$ instead of $\mathcal{D}(I_{pred}),\mathcal{D}(I_{GT})$ to make the loss differentiable [69].
Furthermore, for $p\in\mathcal{D}(I_{pred})$ which are not matched with a point in $\mathcal{D}(I_{GT})$, we match it to the closest point on the diagonal of the persistent diagram, i.e., $b(\eta^{*}(p))$ and $d(\eta^{*}(p))$ equals to 1, 0 if the matched $\eta^{*}(p)$ is a point in $\mathcal{D}(I_{GT})$, otherwise:
$b(\eta^{*}(p))=d(\eta^{*}(p))=\frac{1}{2}(b(p)+d(p))$ (9)
The afore-hand matching process is based on all the candidate matchings, including matching with points in $\mathcal{D}(I_{GT})$ and matching with the diagonal.
Figure 6: Overview of the segmentation network.
III-C3 Loss Formulation
Finally, the persistent topological loss and the perceptual topological loss are used with a conventional pixel-wise BCELoss:
$\mathcal{L}_{bce}=\frac{1}{n}\sum_{i=1}^{n}y_{i}log(x_{i})+(1-y_{i})log(1-x_{i})$ (10)
where $x_{i}$ and $y_{i}$ are the per-pixel predicted likelihood value and the ground truth value. $n$ is the totoal number of pixels in the image. The topological loss functions are controlled by weights $\lambda_{tc}$, $\lambda_{ts}$:
$\mathcal{L}_{total}=\mathcal{L}_{bce}+\lambda_{tc}\mathcal{L}_{tc}+\lambda_{ts}\mathcal{L}_{ts}$ (11)
to jointly optimize the segmentation network.
Figure 7: Different types of masks randomly applied to the retinal images for training both the translation and segmentation networks.
III-C4 Segmentation Network
We use a basic U-Net [28] structure with reduced model size for the segmentation network. In addition, as shown in Fig. 6, we adopt a ‘predict-refine’ strategy which first make a coarse prediction of the vessels and feed the initial prediction and the input image to the network again to refine it. This method has been shown useful in both topology-aware image segmentation [59] and multiple works in deep-learning-based retinal vessel segmentation [38, 39, 40]. In our method, the prediction and refinement networks share the same weight. Additionally, at both the initial prediction stage and the refinement stage, the loss functions are computed to optimize the network.
IV Experiments
TABLE I: Comparison between our UVSM with SOTA finetuned methods. A green box indicates the best result and a blue box indicates the second-best result.
Dataset Modality Camera #img Method Acc. Dice Sp Se Pr F1 MCC clDice $\beta_{0}\downarrow$ $\beta_{1}\downarrow$
$STARE$ $CF$ Topcon 20 W-Net 97.61 81.52 98.80 81.59 82.36 81.52 80.51 85.56 75.6 13.1
WS-DMF 96.01 71.77 97.36 77.28 68.55 71.77 70.30 78.94 193.3 17.3
GT-DLA-dsHff 95.59 66.17 97.70 66.15 66.71 66.17 63.96 66.44 84.0 12.2
U-Net 97.25 78.32 98.67 77.27 80.48 78.32 77.17 81.59 132.7 14.3
CS2-Net 97.80 82.86 98.91 82.85 83.85 82.86 81.97 86.41 84.9 12.8
FA-Net 97.85 83.57 98.84 84.51 83.17 83.57 82.58 87.63 69.1 10.1
Our UVSM 97.48 80.51 98.69 80.77 80.78 80.52 79.31 85.21 35.1 12.6
$ChaseDB1$ $CF$ Zeiss 28 W-Net 97.08 78.87 98.75 76.35 82.32 78.87 77.58 78.46 54.09 9.0
WS-DMF 95.54 59.62 98.81 50.31 78.94 59.62 59.92 58.15 123.6 24.1
GT-DLA-dsHff 96.40 74.26 98.24 72.87 76.11 74.26 72.46 73.06 81.1 7.6
U-Net 96.81 76.84 98.60 74.23 80.36 76.84 75.39 76.01 83.7 8.2
CS2-Net 97.19 80.05 98.65 78.89 81.65 80.05 78.67 80.77 84.9 8.6
FA-Net 97.18 80.21 98.52 80.39 80.72 80.21 78.89 81.29 68.4 7.1
Our UVSM 96.55 77.34 98.05 77.81 75.38 76.34 74.63 78.34 70.1 7.5
$HRF$ $CF$ Canon 35 W-Net 96.60 64.39 98.48 61.70 68.04 64.39 62.88 60.09 260.4 59.8
WS-DMF 94.79 58.02 96.02 71.78 49.75 58.03 56.84 56.54 945.1 51.7
GT-DLA-dsHff 95.49 53.17 97.86 51.16 55.84 53.17 50.99 45.92 421.0 60.0
U-Net 96.58 64.22 98.42 61.82 67.40 64.22 62.64 59.63 297.7 53.1
CS2-Net 96.68 65.17 98.51 62.37 68.82 65.18 63.67 60.82 334.7 51.1
FA-Net 96.67 65.80 98.40 64.36 67.87 65.80 64.23 61.97 257.7 47.5
Our UVSM 96.60 68.50 97.78 74.70 63.87 68.50 67.15 68.26 142.8 55.5
$IOSTAR$ $MC$ Heyex 30 W-Net 97.09 80.66 98.70 78.54 83.50 80.66 79.30 87.46 43.2 9.4
WS-DMF 95.89 71.49 98.29 67.53 77.54 71.49 69.88 76.92 140.9 22.0
GT-DLA-dsHff 96.91 79.63 98.52 78.27 81.67 79.63 78.16 86.99 43.4 7.8
U-Net 97.01 80.50 98.52 79.72 81.80 80.50 79.04 87.08 57.4 8.7
CS2-Net 97.05 80.59 98.59 79.27 82.49 80.59 79.17 87.06 67.2 10.3
FA-Net 97.13 81.38 98.52 81.30 82.23 81.38 80.06 88.08 38.3 10.9
Our UVSM 96.70 78.38 98.36 77.53 79.72 78.38 76.73 85.08 69.1 6.3
$JRCFA$ $FA$ Heyex Optos 40 W-Net 97.83 83.49 99.03 81.73 86.11 83.49 82.57 84.68 162.0 13.7
WS-DMF 97.20 78.20 98.69 76.90 81.84 78.20 77.35 78.60 209.2 14.8
GT-DLA-dsHff 96.38 73.14 98.17 72.25 74.44 73.14 71.31 72.16 279.1 13.2
U-Net 97.89 83.95 99.14 81.65 87.02 83.95 83.04 84.86 160.5 13.9
CS2-Net 97.98 84.88 99.00 84.43 85.79 84.88 83.93 86.07 157.6 11.2
FA-Net 97.94 84.47 99.03 83.51 85.94 84.47 83.51 85.72 126.3 10.1
Our UVSM 96.47 75.00 97.91 78.40 72.89 74.89 73.43 76.37 37.7 15.2
$JRCFAF$ $FAF$ Heyex Optos 40 W-Net 97.93 77.78 99.21 74.01 82.59 77.77 76.98 81.36 48.0 8.1
WS-DMF 97.48 72.19 99.05 67.88 78.94 72.19 71.53 75.63 109.1 10.6
GT-DLA-dsHff 97.24 70.71 98.80 67.71 74.57 70.71 69.50 74.61 62.8 7.5
U-Net 97.89 77.66 99.13 74.94 81.29 77.66 76.81 81.03 68.6 7.2
CS2-Net 97.93 78.11 99.14 75.34 81.64 78.12 77.24 81.46 63.4 6.9
FA-Net 97.95 78.59 99.09 76.69 81.17 78.69 77.70 82.55 42.4 7.3
Our UVSM 97.30 75.49 97.96 85.35 68.19 75.49 74.78 81.89 21.6 4.5
$JRCIR$ IR NIR Heyex 40 W-Net 97.07 79.79 98.85 75.25 83.35 78.80 77.52 79.19 55.3 7.0
WS-DMF 96.32 74.49 98.10 74.37 75.87 74.49 72.88 75.67 107.9 9.8
GT-DLA-dsHff 96.30 73.68 98.30 71.36 76.64 73.68 71.88 74.46 88.5 5.9
U-Net 97.14 79.92 98.66 78.74 81.84 79.92 78.60 80.93 79.4 5.6
CS2-Net 97.22 80.28 98.77 78.30 83.01 80.28 79.01 80.79 86.9 5.6
FA-Net 97.24 80.48 98.75 78.79 82.84 80.48 79.19 81.30 49.6 5.0
Our UVSM 96.10 74.66 97.84 75.31 72.85 73.66 71.81 76.30 28.8 5.5
Figure 8: Qualitative comparison between our UVSM and finetuned SOTA models on multi-modality retinal images. Red, blue orange and green box include the original image, the ground truth label, finetuned FA-Net segmentation and the segmentation from our UVSM. From left to right, up to down: Topcon CF, Canon CF, Heidelberg MC, Optos FA, Heidelberg FA, Optos FAF, Heidelberg FAF, Heidelberg IR.
IV-A Datasets
IV-A1 Image Translation Training
We prepared a multi-modality dataset with over 150k un-identified images from the database of the Jacobs Retina Center at UCSD. However, in our previous numerous attempts we found it very difficult and time-consuming to tune the training of the image translation network (regardless of the methods) for a large dataset. Therefore, in this work, we use a 4k images (2k CF, 500 FA, FAF, MC, IR/NIR, respectively) dataset from the larger 150k database. Each image has a resolution of 768 $\times$ 768. All the CF images are from the 30-degree (field-of-view) Topcon camera and all the IR/NIR images are from the Heidelberg camera. For FA, FAF and MC, 250 images are from Heidelberg camera. For Optos camera, the original images are ultra-wide-filed 135-degree images, which are out of the scope of this work (will be dicussed in section V-A). Therefore, we crop random regions from the center area of the retina to get 30-40-degree conventional field-of-view images and apply a commonly-used circular mask (Fig. 7(a)). In the dataset, there are 250 cropped and masked Optos image for FA, MC, FAF modalities, respectively.
IV-A2 Segmentation Training
We train the segmentation network using only the 20 training images from the public DRIVE [10] CF dataset. In practice, we discard image #14 since the vessel pattern is unusual. In addition, for all the images we resize them to 768$\times$768 for training. We also resized the ground truth annotation and manually corrected the error caused by interpolation. In addition, according to our examination, the original DRIVE ground truth has multiple topological errors such as small redundant holes in the vessel foreground. In our revised ground truth, we correct these errors and allow learning more accurate vessel topological features.
IV-A3 Evaluation
Seven diverse datasets are used for evaluation. For the CF modality, we use the public STARE [78], ChaseDB1 [79] and HRF [16] datasets. For MC modality, we use the public IOSTAR [80] dataset. For FA, FAF and IR/NIR, we prepared three new datasets: JRCFA, JRCFAF and JRCIR. Each dataset has 40 768$\times$768 images, which is more than most of the commonly-used public retinal vessel segmentation (CF) datasets. The JRCFA and JRCFAF dataset consist 20 images from Heidelberg and 20 from Optos, respectively. The JRCIR dataset consists 40 Heidelberg (the only camera for this modality) images. All the vessel annotations are either labeled or carefully examined and refined by the retina experts in the UCSD Jacobs Retina Center (JRC). More information on the 7 datasets can be found in Table I.
IV-B Implementation Details
IV-B1 Translation Network
The balance between the size of the generator and discriminator is important to stably train the CycleGAN for our task. We use a U-Net [28] generator with 16.66M parameters and an NLP discriminator with 2.77M parameters [^2] More details on the network structures we use for translation and segmentation will be provided in our codes.. For training, we use a batch size of 4 and train the model for 100 epochs on the 4000 images. We start with an initial learning rate of $42e-$ 4 and start decaying the learning rate linearly to 0 from epoch 50 to 100. Adam optimizer is used to train both the generators and the discriminators, with a momentum of $(0.5,0.999)$ and no weight decay. $\lambda_{c}=$10.0 and $\lambda_{i}=$0.5.
IV-B2 Segmentation Network
An initial learning rate of 1e-3 is used to train the model for 150 epochs, then decayed to 1e-4 to train for another 250 epochs. Adam optimizer is used with a weight decay of 1e-3 and momentum of $(0.9,0.999)$. We train with a batch size of 4. The weight $\lambda_{tc}$ and $\lambda_{ts}$ we use are 0.05 and 0.0002, respectively. All the testing images are first pad to square and resized to 768$\times$768 to pass through the translation and the segmentation network. Then the prediction is resized back to the original image size to compute the metrics with the ground truth labels. The implementation of both image translation and the segmentation parts are under the PyTorch framework. The computation of persistent homology relies on the GUDHI library [81].
IV-B3 Data Augmentation
Extensive data augmentations are used for training both the translation and the segmentation network. For translation network, random scaling, sharpness, contrast and gamma adjustment are used. Additionally, since different camera manufacturers apply different shapes of mask to the image, we randomly apply the masks shown in Fig. 7 to the images to prevent the model from generating artifacts out of the effective imaging area. For segmentation network, in addition to the aforementioned methods, we also apply random negative film followed by a contrast or sharpness adjustment. This is essential to allow the universal segmentation model to work for the FA modality as the vessels appears to be brighter than the background and we found no image translation method capable of stably inverting the color of the vessels. In addition, random image rotation, shift and Gaussian noise are applied.
Figure 9: Ablation studies on how each proposed step in the pipeline improves the performance of the universal vessel segmentation model.
Figure 10: Qualitative comparison for how each proposed step imporves the segmentation quality (better zoom in to view). From top to bottom: CF, MC, FA, FAF, IR modality.
IV-C Comparison with Finetuned SOTA Methods
To the best of our knowledge, this work is the first universal retinal vessel segmentation model and there is no previous method that we can fairly compare with. Therefore, we compared with state-of-the-art (SOTA) methods which are finetuned individually on each of the 7 datasets in a fully-supervised fashion. Since we use three new datasets and the works in the retinal vessel segmentation literature have inconsistent picks of datasets, dataset splits and inconsistent choices of evaluation metrics, we prepare a new benchmark where all the methods are evaluated in a consistent way for fair comparison. The baseline methods are run using respective recommended training details and in a five-fold experiment that splits the whole dataset by 5 groups. Each fold uses one group for testing and the other four for training. Eventually, all images are used for testing so as to make sure that the average score is computed on the same images with our universal model (which is directly tested on all the images). The baseline methods include U-Net [28], CS2-Net [30], W-Net [39], FA-Net [40], GT-DLA-dsHff [31] and WS-DMF [38]. The training and testing of the baselines are done in original image size except for the HRF dataset. The images in HRF dataset are too large to be trained on our 24GB GPUs. Therefore, we first pad the image and ground truth to square, then downsampling them to 768$\times$768 for training and testing. The predicted vessel likelihood is then upsampled to the original image size to compare with the original ground truth and compute the metrics.
Since the use of evaluation metrics is inconsistent in the literature, in our work we use most of the commonly-used metrics, including the pixel-wise accuracy (Acc.), the Dice score, the specificity (Sp), the sensitivity (Se), the Precision (Pr), the F1 score and the Mathew Correlation Coefficients (MCC). In addition, since topological learning is an important part of our work and the topological accuracy is important in retinal vessel segmentation, we also use three topological metrics which are commonly used in the segmentation of more general tubular structures. These metrics include the clDice score [62] and the 0-/1-Betti errors ($\beta_{0},\beta_{1}$) [63]. All of the metrics are the larger the better except for the 0-/1- Betti errors, which are the smaller the better.
Table I, Fig. 8 show that our proposed model, despite allowing universal segmentation, still possesses comparable performance with SOTA finetuned methods in a general sense. However, for some modalities there still exist non-neglectable performance gap, especially for the FA. As shown by an example in the second row of Fig. 8, the FA modality (particularly from the Optos camera) has very high contrast between the vessel and the background. As a result, more ending vanishing vessels are visible on the image and the universal model often fails to identify them. The reasons are two-fold: (1) The image translation introduces non-avoidable detail loss; (2) The topological loss function makes the model more conservative in predicting a pixel as vessel.
Figure 11: Comparison of the final segmentation performance between using different methods for image translation (reporting pixel-wise accuracy).
Figure 12: Qualitative examples of the translated images using different image translation methods. (best zoom in to view)
IV-D Ablation Studies
IV-D1 Pipeline
In this part, we study how the main components in the proposed pipeline improve the segmentation quality. All the experiments in the ablation studies adopt the same training configurations, data augmentations and network structure, unless otherwise stated. We first run a baseline experiment where no image translation nor topological learning is applied. In another word, we simply train a segmentation model on the DRIVE training set and apply the model to the multi-modality retinal images. As shown by Fig. 9 and Fig. 10, the performance of the model is very poor. After applying the image translation (CycleGAN), the segmentation quality notably improves, especially for the non-CF modalities. However, we still observe some results where the vessel boundary is not clear or adjacent vessels are mixing up (Fig. 10). After further introducing the topological learning, the model makes clean and accurate segmentation.
IV-D2 Image Translation Method
We compare CycleGAN with two most recent Diffusion model-based methods (StyleDiffusion [58] and UNSB [53]). The 19 translated DRIVE images are from the respective translation method to train the downstream segmentation network.
For StyleDiffusion, we randomly pick only one CF image as the style target and reduce the training set size to the same as [58] uses because we find on larger training set StyleDiffusion generates severe artifacts despite our efforts to finetune the training hyperparameters for our images. In addition, data augmentation cannot be implemented in StyleDiffusion for its two-diffusion-stage design. In general, the final segmentation performance by StyleDiffusion is the poorest (Fig. 11). Furthermore, similar with the observation in [12], where style transfer method is also used for retinal vessel segmentation, the transferred images tend to include many vessel-like artifacts (Fig. 12) introduced by the perceptual inductive bias, which will later result in many artifacts in the segmentation.
For UNSB, the training is performed on the same dataset for training CycleGAN. We use the recommended training settings and network structure in [53] except for tuning the number of the diffusion steps to 6 to adapt to our data.
As shown in Fig. 11, when comparing the final segmentation performance, CycleGAN demonstrates the best performance for most datasets. It is also observed that the CycleGAN translation results are visually more alike with the target CF images where the sharpness and contrast of the images are weaker than in the other modalities (Fig. 12). Furthermore, for FAF where the original images are very noisy, CycleGAN also performs better in reducing the noise and generate a smoother translated image.
IV-D3 Topological Segmentation Loss Function
In this part we compare how the combinations of topological segmentation loss functions impact the performance of the UVSM. Table II and Fig. 13 shows that, both $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$ helps with overcoming the domain gap of the non-CF modalities. It is observed that $\mathcal{L}_{tc}$ is unable to address the inverted vessel color in FA images and tend to instead predict the boundary of the vessel, which results in a ‘sandwich’ pattern. In addition, using only $\mathcal{L}_{tc}$ adds saw artifacts and discontinuities to the segmentation. $\mathcal{L}_{ts}$ is able to address the problem of FA modality and improves the discontinuities along the vessel, however, it is observed that if only using $\mathcal{L}_{ts}$, the vessel boundary will be smoothed and the morphological quality of the segmentation will be poor. With a combination of $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$, the saw and smoothing artifacts counteracts, and the challenge in predicting FA vessels is solved, producing the optimal results.
Figure 13: Qualitative results for section IV-D3, From top to bottom: CF, MC, FA, FAF, IR.
TABLE II: Comparison between different choices of topological loss functions. Light blue-conventional pixel-wise metrics, light red-topological metrics, light green-hybrid metric.
Dataset $\mathcal{L}_{tc}$ $\mathcal{L}_{ts}$ Acc. Dice clDice $\beta_{0}$ $\beta_{1}$
$STARE$ 97.28 79.25 83.62 65.5 16.00
$\checkmark$ - 97.29 78.83 83.14 114.7 16.2
- $\checkmark$ 97.08 78.20 82.86 10.7 15.4
$\checkmark$ $\checkmark$ 97.48 80.51 85.21 35.1 12.6
$ChaseDB1$ 96.57 77.03 78.34 92.7 10.5
$\checkmark$ - 96.56 76.62 77.62 151.6 30.8
- $\checkmark$ 96.34 75.32 77.22 19.1 6.4
$\checkmark$ $\checkmark$ 96.55 77.34 78.34 70.1 7.5
$HRF$ 96.52 67.56 66.55 160.5 66.7
$\checkmark$ - 96.59 67.79 66.61 168.5 57.5
- $\checkmark$ 95.99 63.43 61.67 7.4 50.9
$\checkmark$ $\checkmark$ 96.60 68.50 68.26 70.6 55.5
$IOSTAR$ 96.48 77.51 84.61 87.6 8.0
$\checkmark$ - 96.56 77.49 83.56 160.7 6.3
- $\checkmark$ 96.28 76.15 83.01 15.6 6.6
$\checkmark$ $\checkmark$ 96.70 78.38 85.08 69.1 6.3
$JRCFA$ 90.85 40.15 36.35 233.4 17.1
$\checkmark$ - 94.00 58.73 60.22 325.4 98.8
- $\checkmark$ 95.41 69.78 72.91 15.6 16.0
$\checkmark$ $\checkmark$ 96.70 75.00 76.37 37.7 15.2
$JRCFAF$ 95.90 64.45 69.03 203.9 8.9
$\checkmark$ - 97.03 73.46 77.88 149.7 8.8
- $\checkmark$ 96.52 71.00 79.07 17.1 5.6
$\checkmark$ $\checkmark$ 97.30 75.49 81.89 21.6 4.5
$JRCIR$ 95.36 69.77 71.34 144.0 6.7
$\checkmark$ - 96.01 73.48 74.66 149.2 14.6
- $\checkmark$ 95.69 72.09 74.69 21.0 5.4
$\checkmark$ $\checkmark$ 96.10 74.66 76.30 28.8 5.6
V Conclusions
V-A Limitations
As demonstrated and discussed in the experiment section, our current universal segmentation is imperfect and has several limitations to improve upon. Firstly, although the primary vessels are well predicted, the model tends to overprune the vessels, which leads to more false negative pixels and the thin vanishing vessels appearing at the peripheral area of the retina or near the macula are missing in the segmentation. This is caused by both the non-neglectable detail loss of the image translation process and the effect of topological loss function. As a result, performance gap still exists between our method and finetuned SOTA methods. Secondly, our current work only addresses the vessel segmentation problem for multi-modality retinal images in a most commonly-used 30-60 degree field-of-view. While there is an emerging type of 135 degree ultra-wide-field retinal images which our work has not included so far. In addition, there is another 10-15 degree OCTA retinal image which is also commonly used but is not only different in field-of-view, but also in the imaging principle. These two types of images are considerably different with the conventional modalities and requires further research to be included into the universal vessel segmentation model.
V-B Conclusion
In this paper, we proposal a foundational universal vessel segmentation model for multi-modality retinal images. Compared with the past literature of retinal vessel segmentation where (1) at most time only Color Fundus modality is studied and; (2) separate models need to be trained for different datasets (from different cameras) to achieve good performance, our method is able to perform robust vessel segmentation for all commonly-used modalities from all commonly-used cameras using only a single model and achieved comparable performance with SOTA finetuned methods. Although the current model still has limitations and we are working to further improve it, we believe this work is a milestone in the community of retinal image processing and understanding and marks a new track in the research of retinal vessel segmentation.
V-C Future Work
Our future work includes improving the universal vessel segmentation model to allow them to work for a wider range of field of views and to be more robust to images with worse imaging quality. We also aim for systematic study of how to allow the model to be more robust to the outlier patterns which the segmentation training set does not include, especially for severe pathological patterns on the retina (currently we find the topology-aware segmentation method helps addressing this issue, but since we do not have organized dataset for evaluating the problem, we did not report this topic in this paper).

### v2

@@
 However, multi-modality retinal images are used every day in the study of retina and retinal diseases, and the study of vessel segmentation on the other modalities is scarce; (2) Even though a small amount of works extended their experiments to limited new modalities such as the Multi-Color Scanning Laser Ophthalmoscopy (MC), these works still require finetuning a separate model for the new modality.
-And the finetuning will require extra training data, which is difficult to acquire.
+The finetuning will require extra training data, which is difficult to acquire.
 In this work, we present a foundational universal vessel segmentation model (UVSM) for multi-modality retinal images.
-Not only do we perform the study on a much wider range of modalities, but also we propose a universal model to segment the vessels in all these commonly-used modalities.
+Not only do we perform the study on a much wider range of modalities, but we also propose a universal model to segment the vessels in all these commonly-used modalities.
 Despite being much more versatile comparing with existing methods, our universal model still demonstrates comparable performance with the state-of-the-art finetuned methods.
@@
 Index Terms:
-retina, retinal images, vessel segmentation, image translation, topology-preserving image segmentation.
+retina, retinal images, vessel segmentation, image translation, topology-preserving image segmentation, foundation model.
 I Background and Introduction
@@
 Additionally, the pathological information provided by CF is limited.
-Therefore, in more recent years, CF is less and less used and emerging new imaging techniques become increasingly popular.
+Therefore, in more recent years, CF is used less frequently and emerging new imaging techniques become increasingly popular.
 These modalities include (Fig.
-1): (1) Multi-Color Scanning Laser Ophthalmoscopy (MC), which uses more intense laser beam and allow exposing structures in deeper layers of the retina; (2) Florescence Angiography (FA), which uses florescence dye to illuminate the retinal vessels and are commonly used in the diagnosis of diseases related to neovascularization; (3) Fundus AutoFlorescence (FAF), which does not rely on external light source but use natural fluorescence of the retina and provide insights into the health and metabolism of the retinal pigment epithelium (RPE) and photoreceptor cells; (4) Infrared/Near-Infrared Reflectance (IR/NIR), which uses Infrared or Near-Infrared lights to illuminate the image and allows seeing deeper features on the retina such as RPE abnormalities.
-(5) Optical Coherence Tomograph Angiograph (OCTA), which uses low-coherence interferometry to measure the backscattered light from retinal layers.
+1): (1) Multi-Color Scanning Laser Ophthalmoscopy (MC), which uses more intense laser beam and allow exposing structures in deeper layers of the retina; (2) Florescence Angiography (FA), which uses florescence dye to illuminate the retinal vessels and are commonly used in the diagnosis of diseases related to neovascularization; (3) Fundus AutoFlorescence (FAF), which does not rely on external light source but uses natural fluorescence of the retina and provides insights into the health and metabolism of the retinal pigment epithelium (RPE) and photoreceptor cells; (4) Infrared/Near-Infrared Reflectance (IR/NIR), which uses Infrared or Near-Infrared lights to illuminate the image and allows seeing deeper features on the retina such as RPE abnormalities.
+(5) Optical Coherence Tomography Angiography (OCTA), which uses low-coherence interferometry to measure the backscattered light from retinal layers.
 It is primarily used to image the micro-vascular structure on the retina.
@@
 Therefore, in this work we focus on the first four modalities and the CF.
-Figure 1: Overview of multi-modality retinal images and the results from our UVSM.
+Figure 1: Overview of multi-modality retinal images and the segmentation results from our UVSM.
 In each image, the upper half is the original image and the lower half is the segmentation.
@@
 In this work, we prepare three high-quality datasets for FA, FAF and IR, respectively.
-Each dataset has 40 images and the vessels are professionally and carefully labeled.
+Each dataset has 40 images and the vessels are carefully labeled by retina experts.
 Nevertheless, despite the availability of the new datasets, in this work, we only use them for evaluation purpose.
@@
 There are limited previous works which are extended beyond CF modality [11, 12, 13].
-However, not only they did not work with as comprehensive modalities as we do, but also in these work one still needs to finetune a separate model on the target modality and each model only works for one modality.
-We instead aim to build a universal model to use a single model and allows robust segmentation for all the commonly-used modalities.
+In addition to lacking comprehensive modalities (compared to our approach), one still needs to finetune a separate model on the target modality where each model only works for one image modality.
+We instead aim to build a single universal model that achieves robust segmentation for all the commonly-used modalities.
 The model will serve as an important tool in cross-modality retinal image study, which becomes increasingly important in modern research of the retina.
-To achieve the universal segmentation with limited labeled data.
-We propose our model in two major steps (Fig.
+To achieve the universal segmentation with limited labeled data, we develop our model in two major steps (Fig.
 2).
@@
 The topology-aware segmentation methods were initially proposed to improve the topological accuracy of the segmentation of tubular structures within a certain type of image.
-We extend these methods to our task and found that not only do they improve the topological accuracy in the segmentation (e.g., preserving vessel continuity, branching and loop features), but also they force the segmentation to be topologically correct so when images with slight domain gap are inputted, the model can still make the correct segmentation.
+We extend these methods to our task and found that they improve the topological accuracy in the segmentation (e.g., preserving vessel continuity, branching and loop features), as well as they force the segmentation to be topologically correct so when images with slight domain gap are inputted, the model can still make the correct segmentation.
 Our universal vessel segmentation model is evaluated on 7 diverse datasets (including 3 new datasets we prepared for this work) of the 5 most commonly-used retinal image modalities (CF, FA, FAF, MC, IR/NIR) taken from 5 most commonly-used cameras (Topcon, Canon, Zeiss, Heidelberg Spectralis, Optos).
@@
 Extensive ablation studies were also performed to investigate how different design choices in the proposed pipeline affects the performance of the universal vessel segmentation.
-The main contributions of this paper can be summerized as:
+The main contributions of this paper are:
   • We propose a foundational universal vessel segmentation model (UVSM) for multi-modality retinal images which can perform robust vessel segmentation for all the commonly-used modalities and cameras using one single model without providing any modality information.
@@
 II-A Retinal Vessel Segmentation
-Retinal vessel segmentation classifies each pixel on the retinal images into foreground (vessel) and background (non-vessel).
-And the literature focuses on segmenting vessels in the CF images.
-Various machine learning-based or non-machine learning-based methods were proposed.
-A majority of the early methods are based on the matched filter [14, 15, 16], which convolves retinal images with predefined or learned kernels and then adaptively threshold the image to obtain the vessel segmentation map.
-Another type of methods use tracking algorithms to mathematically estimate the growth of the retinal vessel trees from seed points [17, 18].
+Retinal vessel segmentation classifies each pixel in the retinal images into foreground (vessel) and background (non-vessel).
+Current works focus on segmenting vessels in the CF images, where various machine learning-based or non-machine learning-based methods were proposed.
+A majority of the early methods are based on the matched filter [14, 15, 16], which convolve retinal images with predefined or learned kernels and then adaptively threshold the image to obtain the vessel segmentation map.
+Another type of methods uses tracking algorithms to mathematically estimate the growth of the retinal vessel trees from seed points [17, 18].
 The segmentation can then be derived by extending around the estimated vessel centerlines using the estimated vessel diameters.
 Some methods also adopt morphological image processing methods such as the top-hat operation to extract the vessel or as a post-processing step [19, 20].
-Furthermore, numerous traditional machine learning-based methods were also applied in retinal vessel segmentation.
-Many classic models were used, such as the support vector machine [21], random forest [22] and Adaboost [23] for supervised methods, and Gaussian mixture model [24], Fuzzy C-Means [25] for unsupervised methods.
+Furthermore, numerous traditional machine learning-based methods were also applied in retinal vessel segmentation such as the support vector machine [21], random forest [22] and Adaboost [23] for supervised methods, and Gaussian mixture model [24], Fuzzy C-Means [25] for unsupervised methods.
 In more recent years, deep learning-based methods dominate the literature.
-They majorly rely on the Convolutional Neural Networks (CNN) as a backbone model due to the existence of vast fine details and the whole-spreading characteristic of the vessels.
-An introductory work is the Deep Retinal Image Understanding (DRIU) [26], which adopts a VGG [27] network structure as encoder and assign specialized output layers to the embedded features to segment the vessels and the optic disc from the retina.
+They mainly rely on the Convolutional Neural Networks (CNN) as a backbone model due to the existence of vast fine details and the spread-out characteristic of the vessels.
+An introductory work is the Deep Retinal Image Understanding (DRIU) [26], which adopts a VGG [27] network structure as encoder and assigns specialized output layers to the embedded features to segment the vessels and the optic disc from the retina.
 Based on a widely-used U-Net [28] network structure and to seek further improvement for retinal vessel segmentation, [29, 30, 31, 32, 33] proposed to add various global reasoning modules at the bottom of the network.
@@
 Topology-aware image segmentation aims to preserve the topological accuracy in the segmentation of tubular and net-like structures.
-The literature mainly researches a topological loss function that extracts topological features from the ground truth and segmentation and minimizes the difference between them.
-One type of methods extract certain features from the segmentation which are proximity to the actual topological features of the foreground objects.
+The literature mainly considers a topological loss function that extracts topological features from the ground truth and segmentation and minimizes the difference between them.
+One type of methods extracts certain features from the segmentation which are proximity to the actual topological features of the foreground objects.
 A pioneer work [59] uses an ImageNet-pretrained VGG [60, 27] network after the segmentation network.
@@
 The topological loss is then computed between the ground truth and prediction perceptual features.
-Another topological segmentation loss function [61] was proposed exclusively for the the segmentation of glands.
+Another topological segmentation loss function [61] was proposed exclusively for the segmentation of glands.
 The loss is computed according to the iterations needed to derive the gland skeletons by the image erosion operation.
@@
 Using such a relationship as a loss function is shown to strengthen the connectivity along the tubular structure thus providing better topological accuracy.
-Another type of methods use strict topological features of the segmentation which are based on the computational topology theory and receieved more research attentions in recent years.
+Another type of methods uses strict topological features of the segmentation which are based on the computational topology theory and received more research attentions in recent years.
 The topological features are mainly based on the persistent homology theory [63, 64, 65, 66, 37, 67, 68, 69].
-A revolutionary work that first introduce this concept in image segmenation is [63], which computes the persistent features from the ground truth and prediction, matching the features using a Wasserstein distance [70] and use the distance as the loss function.
+A revolutionary work that first introduce this concept in image segmenation is [63], which computes the persistent features from the ground truth and prediction, matching the features using a Wasserstein distance [70] and using the distance as the loss function.
 The following works share similar idea.
@@
 [67] proposed to match the longest persistent barcodes in the prediction to the ground truth.
-And [37] proposed to use a Hausdorff distance [71] to matches the persistent diagrams to improve the sensitivity of the matching to the outliers.
-[69] and [68] both aim to compute more precise persistent feature matching, as in earlier methods the actual matchings results are highly inaccurate due to only considering the global topological features.
+[37] proposed to use a Hausdorff distance [71] to match the persistent diagrams to improve the sensitivity of the matching to the outliers.
+[69] and [68] both aim to compute more precise persistent feature matching, as in earlier methods the actual matching results are highly inaccurate due to only considering the global topological features.
 [69] and [68] instead allow the loss to act more selectively on the topological features thus improve the efficiency of the topological loss function.
 [68] proposed to use an Induced Matching [72] method which constructs a common ambient space between the ground truth and prediction and maps the persistent features in between to find a better matching.
-However, the computational cost of this method is very high, especially for large images.
-And the retinal images usually have high resolutions.
+However, the computational cost of this method is very high, especially for large images, where typical retinal images usually have high resolutions.
 Alternatively, [69] proposed to leverage the spatial locations of the persistent features referencing their critical cell locations for more faithful persistent feature matching and is computationally much more efficient than [68].
@@
 Therefore, we pick CycleGAN as the backbone of our image translation model.
-Consequently, the following sections will be discussing the implementation based on CycleGAN and we refer the readers to [58, 53] for details about the other two methods and to section IV-D2 for our implementation details.
+Consequently, the following sections discuss the implementation based on CycleGAN and we refer the readers to [58, 53] for details about the other two methods and to section IV-D2 for our implementation details.
 Figure 3: Overview of our image translation network, which is based on CycleGAN [41].
@@
 3.
-To be more specific, we first update the generators for both $I_{A}$ and $I_{B}$, and then update the discriminators.
+In addition, we first update the generators for both $I_{A}$ and $I_{B}$, and then update the discriminators.
 III-C Topology-aware Segmentation Network
@@
 $[\inf_{\eta:\mathcal{D}(I_{pred})\rightarrow\mathcal{D}(I_{GT})}\sum_{p\in\mathcal{D}(I_{pred})}s_{p}\left\|p-\eta(p))\right\|^{q}]^{\frac{1}{q}}$
-where $s_{p}=\left\|c_{b}(p)-c_{b}(\eta(p))\right\|^{q}$ is a spatial weight that is computed between the creation cells of the persistent features.
+where $s_{p}=\left\|c_{b}(p)-c_{b}(\eta(p))\right\|^{q}$ is a spatial weight that is computed between the creation cells ($c_{b}(\cdot)$) of the persistent features.
 $p$ and $\eta(p)$ are a persistent feature from $I_{pred}$ and a candidate persistent feature from $I_{GT}$ to be matched with.
@@
 where $x_{i}$ and $y_{i}$ are the per-pixel predicted likelihood value and the ground truth value.
-$n$ is the totoal number of pixels in the image.
+$n$ is the total number of pixels in the image.
 The topological loss functions are controlled by weights $\lambda_{tc}$, $\lambda_{ts}$:
@@
 IV Experiments
-TABLE I: Comparison between our UVSM with SOTA finetuned methods. A green box indicates the best result and a blue box indicates the second-best result.
+TABLE I: Comparison between our UVSM with SOTA finetuned methods. A green box indicates the best result and a blue box indicates the second-best result. Heidel refers to Heidelberg Spectralis.
 Dataset Modality Camera #img Method Acc. Dice Sp Se Pr F1 MCC clDice $\beta_{0}\downarrow$ $\beta_{1}\downarrow$
@@
 Our UVSM 96.60 68.50 97.78 74.70 63.87 68.50 67.15 68.26 142.8 55.5
-$IOSTAR$ $MC$ Heyex 30 W-Net 97.09 80.66 98.70 78.54 83.50 80.66 79.30 87.46 43.2 9.4
+$IOSTAR$ $MC$ Heidel 30 W-Net 97.09 80.66 98.70 78.54 83.50 80.66 79.30 87.46 43.2 9.4
 WS-DMF 95.89 71.49 98.29 67.53 77.54 71.49 69.88 76.92 140.9 22.0
@@
 Our UVSM 96.70 78.38 98.36 77.53 79.72 78.38 76.73 85.08 69.1 6.3
-$JRCFA$ $FA$ Heyex Optos 40 W-Net 97.83 83.49 99.03 81.73 86.11 83.49 82.57 84.68 162.0 13.7
+$JRCFA$ $FA$ Heidel Optos 40 W-Net 97.83 83.49 99.03 81.73 86.11 83.49 82.57 84.68 162.0 13.7
 WS-DMF 97.20 78.20 98.69 76.90 81.84 78.20 77.35 78.60 209.2 14.8
@@
 Our UVSM 96.47 75.00 97.91 78.40 72.89 74.89 73.43 76.37 37.7 15.2
-$JRCFAF$ $FAF$ Heyex Optos 40 W-Net 97.93 77.78 99.21 74.01 82.59 77.77 76.98 81.36 48.0 8.1
+$JRCFAF$ $FAF$ Heidel Optos 40 W-Net 97.93 77.78 99.21 74.01 82.59 77.77 76.98 81.36 48.0 8.1
 WS-DMF 97.48 72.19 99.05 67.88 78.94 72.19 71.53 75.63 109.1 10.6
@@
 Our UVSM 97.30 75.49 97.96 85.35 68.19 75.49 74.78 81.89 21.6 4.5
-$JRCIR$ IR NIR Heyex 40 W-Net 97.07 79.79 98.85 75.25 83.35 78.80 77.52 79.19 55.3 7.0
+$JRCIR$ IR NIR Heidel 40 W-Net 97.07 79.79 98.85 75.25 83.35 78.80 77.52 79.19 55.3 7.0
 WS-DMF 96.32 74.49 98.10 74.37 75.87 74.49 72.88 75.67 107.9 9.8
@@
 Our UVSM 96.10 74.66 97.84 75.31 72.85 73.66 71.81 76.30 28.8 5.5
-Figure 8: Qualitative comparison between our UVSM and finetuned SOTA models on multi-modality retinal images.
+Figure 8: Qualitative comparison between our UVSM and finetuned SOTA models on multi-modality retinal images (better zoom in to view).
 Red, blue orange and green box include the original image, the ground truth label, finetuned FA-Net segmentation and the segmentation from our UVSM.
-From left to right, up to down: Topcon CF, Canon CF, Heidelberg MC, Optos FA, Heidelberg FA, Optos FAF, Heidelberg FAF, Heidelberg IR.
+From left to right, up to down: Topcon CF, Canon CF, Zeiss CF, Heidelberg MC, Optos FA, Heidelberg FA, Optos FAF, Heidelberg FAF, Heidelberg IR.
+Bottom right shows the qualitative results for Optos MC and Canon FA (no ground truth available to compare with and to train the baseline finetuned models).
 IV-A Datasets
 IV-A1 Image Translation Training
-We prepared a multi-modality dataset with over 150k un-identified images from the database of the Jacobs Retina Center at UCSD.
+We prepared a multi-modality database with over 150k de-identified images from the database of the Jacobs Retina Center at UCSD.
 However, in our previous numerous attempts we found it very difficult and time-consuming to tune the training of the image translation network (regardless of the methods) for a large dataset.
@@
 Each dataset has 40 768$\times$768 images, which is more than most of the commonly-used public retinal vessel segmentation (CF) datasets.
-The JRCFA and JRCFAF dataset consist 20 images from Heidelberg and 20 from Optos, respectively.
+The JRCFA and JRCFAF datasets both consist 20 images from Heidelberg and 20 from Optos, respectively.
 The JRCIR dataset consists 40 Heidelberg (the only camera for this modality) images.
@@
 More information on the 7 datasets can be found in Table I.
+In addition to the seven vessel segmentation dataset, we also evaluate our model qualitatively on the Optos MC images using images from our 150K database and on the Canon FA images on the CF-FA registration dataset [81].
+We provide preliminary insights in this paper (Fig.
+1, Fig.
+8) yet full evaluation requires the vessel labels (unavailable) and is left as future work.
 IV-B Implementation Details
@@
 For training, we use a batch size of 4 and train the model for 100 epochs on the 4000 images.
-We start with an initial learning rate of $42e-$ 4 and start decaying the learning rate linearly to 0 from epoch 50 to 100.
+We start with an initial learning rate of 2e-4 and start decaying the learning rate linearly to 0 from epoch 50 to 100.
 Adam optimizer is used to train both the generators and the discriminators, with a momentum of $(0.5,0.999)$ and no weight decay.
@@
 We train with a batch size of 4.
-The weight $\lambda_{tc}$ and $\lambda_{ts}$ we use are 0.05 and 0.0002, respectively.
-All the testing images are first pad to square and resized to 768$\times$768 to pass through the translation and the segmentation network.
+The weights $\lambda_{tc}$ and $\lambda_{ts}$ are 0.05 and 0.0002, respectively.
+All the testing images are first padded to square and resized to 768$\times$768 to pass through the translation and the segmentation network.
 Then the prediction is resized back to the original image size to compute the metrics with the ground truth labels.
@@
 Therefore, we compared with state-of-the-art (SOTA) methods which are finetuned individually on each of the 7 datasets in a fully-supervised fashion.
-Since we use three new datasets and the works in the retinal vessel segmentation literature have inconsistent picks of datasets, dataset splits and inconsistent choices of evaluation metrics, we prepare a new benchmark where all the methods are evaluated in a consistent way for fair comparison.
+Since we use three new datasets and the works in the retinal vessel segmentation literature have inconsistent dataset choice, dataset split, and inconsistent choice of evaluation metrics, we prepare a new benchmark where all the methods are evaluated in a consistent way for fair comparison.
 The baseline methods are run using respective recommended training details and in a five-fold experiment that splits the whole dataset by 5 groups.
@@
 All of the metrics are the larger the better except for the 0-/1- Betti errors, which are the smaller the better.
-Table I, Fig.
+For all of them reported metrics, we compute them at the original image size and averaged over all test images.
+The results in Table I, Fig.
 8 show that our proposed model, despite allowing universal segmentation, still possesses comparable performance with SOTA finetuned methods in a general sense.
@@
 The 19 translated DRIVE images are from the respective translation method to train the downstream segmentation network.
-For StyleDiffusion, we randomly pick only one CF image as the style target and reduce the training set size to the same as [58] uses because we find on larger training set StyleDiffusion generates severe artifacts despite our efforts to finetune the training hyperparameters for our images.
+For StyleDiffusion, we randomly pick only one CF image as the style target and reduce the training set size to the same as that in [58] because we find that on larger training set, StyleDiffusion generates severe artifacts despite our efforts to finetune the training hyperparameters for our images.
 In addition, data augmentation cannot be implemented in StyleDiffusion for its two-diffusion-stage design.
@@
 12).
-Furthermore, for FAF where the original images are very noisy, CycleGAN also performs better in reducing the noise and generate a smoother translated image.
-IV-D3 Topological Segmentation Loss Function
-In this part we compare how the combinations of topological segmentation loss functions impact the performance of the UVSM.
-Table II and Fig.
-13 shows that, both $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$ helps with overcoming the domain gap of the non-CF modalities.
-It is observed that $\mathcal{L}_{tc}$ is unable to address the inverted vessel color in FA images and tend to instead predict the boundary of the vessel, which results in a ‘sandwich’ pattern.
-In addition, using only $\mathcal{L}_{tc}$ adds saw artifacts and discontinuities to the segmentation.
-$\mathcal{L}_{ts}$ is able to address the problem of FA modality and improves the discontinuities along the vessel, however, it is observed that if only using $\mathcal{L}_{ts}$, the vessel boundary will be smoothed and the morphological quality of the segmentation will be poor.
-With a combination of $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$, the saw and smoothing artifacts counteracts, and the challenge in predicting FA vessels is solved, producing the optimal results.
+Furthermore, for FAF where the original images are very noisy, CycleGAN also performs better in reducing the noise and generating a smoother translated image.
 Figure 13: Qualitative results for section IV-D3, From top to bottom: CF, MC, FA, FAF, IR.
@@
 $\checkmark$ $\checkmark$ 96.10 74.66 76.30 28.8 5.6
+IV-D3 Topological Segmentation Loss Function
+In this part we compare how the combinations of topological segmentation loss functions impact the performance of the UVSM.
+Table II and Fig.
+13 show that, both $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$ helps with overcoming the domain gap of the non-CF modalities.
+It is observed that $\mathcal{L}_{tc}$ is unable to address the inverted vessel color in FA images and tend to instead predict the boundary of the vessel, which results in a ‘sandwich’ pattern.
+In addition, using only $\mathcal{L}_{tc}$ adds saw artifacts and discontinuities to the segmentation.
+$\mathcal{L}_{ts}$ is able to address the problem of FA modality and improves the discontinuities along the vessel, however, it is observed that if only $\mathcal{L}_{ts}$ is used, the vessel boundary will be smoothed and the morphological quality of the segmentation will be poor.
+With a combination of $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$, the saw and smoothing artifacts counteracts, and the challenge in predicting FA vessels is solved, producing the optimal results.
 V Conclusions
 V-A Limitations
-As demonstrated and discussed in the experiment section, our current universal segmentation is imperfect and has several limitations to improve upon.
-Firstly, although the primary vessels are well predicted, the model tends to overprune the vessels, which leads to more false negative pixels and the thin vanishing vessels appearing at the peripheral area of the retina or near the macula are missing in the segmentation.
+As demonstrated and discussed in the experiment section, our current universal segmentation has several limitations to improve upon.
+Firstly, although the primary vessels are well predicted, the model tends to overprune the vessels, which leads to thin vanishing vessels appearing at the peripheral area of the retina or near the macula are missing in the segmentation.
 This is caused by both the non-neglectable detail loss of the image translation process and the effect of topological loss function.
-As a result, performance gap still exists between our method and finetuned SOTA methods.
 Secondly, our current work only addresses the vessel segmentation problem for multi-modality retinal images in a most commonly-used 30-60 degree field-of-view.
-While there is an emerging type of 135 degree ultra-wide-field retinal images which our work has not included so far.
-In addition, there is another 10-15 degree OCTA retinal image which is also commonly used but is not only different in field-of-view, but also in the imaging principle.
+Our work does not include the ultra-wide-field 135 degrees retina images.
+Moreover, we does not consider the commonly-used 10-15 degree OCTA image, as its imaging principle is different.
 These two types of images are considerably different with the conventional modalities and requires further research to be included into the universal vessel segmentation model.
-V-B Conclusion
-In this paper, we proposal a foundational universal vessel segmentation model for multi-modality retinal images.
+V-B Conclusion and Future Work
+In this paper, we propose a foundational universal vessel segmentation model for multi-modality retinal images.
 Compared with the past literature of retinal vessel segmentation where (1) at most time only Color Fundus modality is studied and; (2) separate models need to be trained for different datasets (from different cameras) to achieve good performance, our method is able to perform robust vessel segmentation for all commonly-used modalities from all commonly-used cameras using only a single model and achieved comparable performance with SOTA finetuned methods.
-Although the current model still has limitations and we are working to further improve it, we believe this work is a milestone in the community of retinal image processing and understanding and marks a new track in the research of retinal vessel segmentation.
-V-C Future Work
-Our future work includes improving the universal vessel segmentation model to allow them to work for a wider range of field of views and to be more robust to images with worse imaging quality.
-We also aim for systematic study of how to allow the model to be more robust to the outlier patterns which the segmentation training set does not include, especially for severe pathological patterns on the retina (currently we find the topology-aware segmentation method helps addressing this issue, but since we do not have organized dataset for evaluating the problem, we did not report this topic in this paper).
+Although the current model still has limitations and we are working to further improve it, we believe this work is a milestone in the community of retinal image processing and understanding and marks a new direction in the research of retinal vessel segmentation.
+Our future work includes improving the universal vessel segmentation model to allow them to work for a wider range of field of views and to better predict the fine vanishing vessels.

### v3

@@
 We identify two major limitations in the existing studies on retinal vessel segmentation: (1) Most existing works are restricted to one modality, i.e, the Color Fundus (CF).
-However, multi-modality retinal images are used every day in the study of retina and retinal diseases, and the study of vessel segmentation on the other modalities is scarce; (2) Even though a small amount of works extended their experiments to limited new modalities such as the Multi-Color Scanning Laser Ophthalmoscopy (MC), these works still require finetuning a separate model for the new modality.
+However, multi-modality retinal images are used every day in the study of retina and diagnosis of retinal diseases, and the study of vessel segmentation on the other modalities is scarce; (2) Even though a few works extended their experiments to limited new modalities such as the Multi-Color Scanning Laser Ophthalmoscopy (MC), these works still require finetuning a separate model for the new modality.
 The finetuning will require extra training data, which is difficult to acquire.
-In this work, we present a foundational universal vessel segmentation model (UVSM) for multi-modality retinal images.
+In this work, we present a novel universal vessel segmentation model (UVSM) for multi-modality retinal images.
 Not only do we perform the study on a much wider range of modalities, but we also propose a universal model to segment the vessels in all these commonly-used modalities.
 Despite being much more versatile comparing with existing methods, our universal model still demonstrates comparable performance with the state-of-the-art finetuned methods.
-To the best of our knowledge, this is the first work that achieves cross-modality retinal vessel segmentation and also the first work to study retinal vessel segmentation in some novel modalities.
+To the best of our knowledge, this is the first work that achieves modality-agnostic retinal vessel segmentation and also the first work that studies retinal vessel segmentation in some novel modalities.
 [^1] Our codes and model will be available at the time of publication.
@@
 Index Terms:
-retina, retinal images, vessel segmentation, image translation, topology-preserving image segmentation, foundation model.
+retina, retinal image processing, vessel segmentation, image translation, topological learning.
 I Background and Introduction
@@
 CF uses white visible light to illuminate the retina and the reflected light is used for generating the image.
-Nevertheless, although CF was the most widely-used modality in the past, it requires dilation before the imaging and is uncomfortable for the patients.
+Nevertheless, although CF was the most widely-used modality in the past, it requires dilation before imaging and is uncomfortable for the patients.
 Additionally, the pathological information provided by CF is limited.
@@
 (5) Optical Coherence Tomography Angiography (OCTA), which uses low-coherence interferometry to measure the backscattered light from retinal layers.
-It is primarily used to image the micro-vascular structure on the retina.
-Nevertheless, OCTA has a much smaller imaging area than the other modalities and has significantly different retinal vessel morphology and topology.
-Therefore, in this work we focus on the first four modalities and the CF.
+OCTA has a much smaller imaging area than the other modalities and has considerably different retinal vessel morphology and topology.
 Figure 1: Overview of multi-modality retinal images and the segmentation results from our UVSM.
@@
 We instead aim to build a single universal model that achieves robust segmentation for all the commonly-used modalities.
-The model will serve as an important tool in cross-modality retinal image study, which becomes increasingly important in modern research of the retina.
+The model could serve as an important tool in cross-modality retinal image study, which becomes increasingly important in modern research of the retina.
 To achieve the universal segmentation with limited labeled data, we develop our model in two major steps (Fig.
@@
 Secondly, we train a segmentation model on the Topcon CF domain to perform the segmentation.
-However, image translation can not fully eliminate the domain gaps between different modalities given such a difficult many-to-one translation task.
+However, image translation cannot fully eliminate the domain gaps between different modalities given such difficult many-to-one translation task.
 To solve this problem, we propose to use a topology-aware segmentation method that learns topological features of the vessels in addition to their conventional pixel-level features.
-The topology-aware segmentation methods were initially proposed to improve the topological accuracy of the segmentation of tubular structures within a certain type of image.
-We extend these methods to our task and found that they improve the topological accuracy in the segmentation (e.g., preserving vessel continuity, branching and loop features), as well as they force the segmentation to be topologically correct so when images with slight domain gap are inputted, the model can still make the correct segmentation.
-Our universal vessel segmentation model is evaluated on 7 diverse datasets (including 3 new datasets we prepared for this work) of the 5 most commonly-used retinal image modalities (CF, FA, FAF, MC, IR/NIR) taken from 5 most commonly-used cameras (Topcon, Canon, Zeiss, Heidelberg Spectralis, Optos).
+The topology-aware segmentation methods were initially proposed to improve the topological accuracy of the segmentation of curvilinear structures.
+We go beyond their conventional applications and propose the topological learning as a novel domain adaptation method for retinal vessel segmentation, in a way that forces the segmentation to be topologically correct so when images from different domains are inputted, the model can still output robust segmentation.
+Our universal vessel segmentation model is evaluated on 10 diverse datasets (including 3 new datasets we prepared for this work) of the 6 most commonly-used retinal image modalities (CF, FA, FAF, MC, IR/NIR, OCTA) taken from 6 most commonly-used camera systems (Topcon, Canon, Zeiss, Heidelberg Spectralis, Optos, Optovue).
 Our universal model is compared with state-of-the-art methods which are finetuned on each of the 7 datasets, respectively, and is shown to achieve comparable performance.
-Extensive ablation studies were also performed to investigate how different design choices in the proposed pipeline affects the performance of the universal vessel segmentation.
+Extensive ablation studies were also performed to investigate how different components in the proposed pipeline affects the performance of the universal vessel segmentation.
 The main contributions of this paper are:
-  • We propose a foundational universal vessel segmentation model (UVSM) for multi-modality retinal images which can perform robust vessel segmentation for all the commonly-used modalities and cameras using one single model without providing any modality information.
+  • We propose the universal retinal vessel segmentation task.
+To the best of our knowledge, this is the first work that treat retinal vessel segmentation as a modality-agnostic segmentation task covering multiple imaging modalities and camera systems, reflecting the real-world need for generalizable retinal vessel segmentation solutions.
+  • We propose the first universal (retinal) vessel segmentation model (UVSM) for multi-modality retinal images based on image translation and image topological learning.
   • We investigate image-to-image translation between retinal images from different modalities.
 We explore different Generative Adversarial Nets (GAN)-based and Diffusion Model-based image translation methods and evaluate their performance on retinal images.
-  • We propose to use a topology-aware image segmentation method to train our segmentation model to allow more robust segmentation across modalities.
-Furthermore, we discover an additional merit of the topology-aware segmentation loss functions in segmenting images with domain gaps in addition to their conventional applications.
+  • We propose topological feature learning as a novel domain adaptation strategy for multi-modal retinal vessel segmentation.
+We show that by learning the topological features of the vessels, the segmentation model becomes much more robust to inputs from different domains (modalities and imaging systems).
   • We collect and annotate three high-quality vessel segmentation datasets for the FA, FAF and IR retinal images, respectively.
-  • We provide a refined version of the widely-used DRIVE dataset with better topological accuracy in the vessel labels (to be discussed in section IV-A2).
+  • We provide a refined version of the widely-used DRIVE dataset with improved topological accuracy in the annotation (to be discussed in section IV-A2).
 II Related Work
 II-A Retinal Vessel Segmentation
-Retinal vessel segmentation classifies each pixel in the retinal images into foreground (vessel) and background (non-vessel).
-Current works focus on segmenting vessels in the CF images, where various machine learning-based or non-machine learning-based methods were proposed.
+Retinal vessel segmentation is a specialized task that general-purpose foundational segmentation models [14, 15, 16] struggle to handle effectively.
+Current works in retinal vessel segmentation focus on segmenting vessels in Color Fundus retinal images, where various machine learning-based or non-machine learning-based methods were proposed.
 A majority of the early methods are based on the matched filter [17, 18, 19], which convolve retinal images with predefined or learned kernels and then adaptively threshold the image to obtain the vessel segmentation map.
@@
 The topological features are mainly based on the persistent homology theory [66, 67, 68, 69, 40, 70, 71, 72].
-A revolutionary work that first introduce this concept in image segmenation is [63], which computes the persistent features from the ground truth and prediction, matching the features using a Wasserstein distance [70] and using the distance as the loss function.
+A revolutionary work that first introduces this concept in image segmenation is [66], which computes the persistent features from the ground truth and prediction, matching the features using a Wasserstein distance [73] and using the distance as the loss function.
 The following works share similar idea.
@@
 All of the training and inference are then performed on the green-channel images.
-Following the pre-processing, we train an domain adaptation model for multi-modality retinal images.
-Although several domain adaptation methods were proposed for retinal vessel segmentation, such as using the pseudo-label methods [39], adversarial learning [33] and asymmetrical maximum classifier discrepancy [73], they were proposed only to address the domain gap between different datasets (from different cameras) in CF images.
+Following pre-processing, we train a domain adaptation model for multi-modality retinal images.
+Although several domain adaptation methods were proposed for retinal vessel segmentation, such as using the pseudo-label methods [42, 76], teacher-student network [77], adversarial learning [36, 76] and asymmetrical maximum classifier discrepancy [78], they were proposed only to address the domain gap between different datasets (from different cameras) in CF images.
 Moreover, these methods still require unsupervised finetuning on the target domain images, which deviate from our goal of universal segmentation.
-Therefore, in this work we propose to use image translation as the domain adaptation method, which allows us to train a single model for all the modalities.
+Therefore, in this work we propose to use image translation as the preliminary domain adaptation method, which allows us to train a single model for all the modalities.
 The image translation model aims to translate images from all the modalities to a uniform Topcon CF modality, and is trained between equivalent number of non-CF images (FA, FAF, MC, IR/NIR) and the Topcon CF images, i.e., during the training, we treat all the non-CF modalities as one type of image.
@@
 Nevertheless, since we are training a difficult many-to-one (FA, FAF, MC, IR/NIR to CF) translation, the translation model cannot perfectly address the domain gap between the non-CF images and the CF images (Fig.
-12).
-We therefore propose to further solve this problem by introducing a topology-aware segmentation method.
-In addition to a conventional pixel-wise loss function (BCELoss), the segmentation network is jointly trained with two topological loss functions.
+13).
+To solve this problem, we propose topology-aware segmentation as a novel domain adaptation method.
+We show that by learning additional topological features of the vessels, the segmentation model becomes much more robust to inputs from different domains.
+In practice, in addition to a conventional pixel-wise loss function (BCELoss), the segmentation network is jointly trained with two topological loss functions.
 At inference, we can disregard the modality of the image.
@@
 For GAN-based method, we experiment with CycleGAN [44].
-We will show in section IV-D2 that CycleGAN, despite being a classic method, outperforms the other two most recent methods in most datasets.
+We will show in section IV-E2 that CycleGAN, despite being a classic method, outperforms the other two most recent methods in most datasets.
 Therefore, we pick CycleGAN as the backbone of our image translation model.
-Consequently, the following sections discuss the implementation based on CycleGAN and we refer the readers to [58, 53] for details about the other two methods and to section IV-D2 for our implementation details.
+Consequently, the following sections discuss the implementation based on CycleGAN and we refer the readers to [61, 56] for details about the other two methods and to section IV-E2 for our implementation details.
 Figure 3: Overview of our image translation network, which is based on CycleGAN [44].
@@
 To train the translation network, we treat all the non-CF images (FA, FAF, MC, IR/NIR) as image type A and the Topcon CF images as image type B.
-At each iteration, a pair of image A batch and image B batch is drawn from the training set.
+At each iteration, a pair of image A batch and image B batch is selected from the training set.
 As illustrated by Fig.
@@
 where the adversarial loss defined as:
-$\textit{L}_{adv.-G_{A}}=\left|D_{A}(G_{A}(I_{A}))-1\right|^{2}$ (2)
+$\textit{L}_{adv.-G_{A}}=\left|D_{A}(G_{A}(I_{A}))-1\right|^{2}_{2}$ (2)
 Furthermore, the translated fake CF images are again fed into the opposite generator to reconstruct the non-CF images. The reconstruction is supervised by a cycle consistency loss:
-$\textit{L}_{cycle-A}=\left|I_{A}-G_{B}(G_{A}(I_{A}))\right|$ (3)
+$\textit{L}_{cycle-A}=\left|I_{A}-G_{B}(G_{A}(I_{A}))\right|_{2}$ (3)
 and is controlled by a weight $\lambda_{c}$. Finally, the real images are also passed through their opposite generator to generate the image itself. This step is supervised with an identity loss:
-$\textit{L}_{identity-A}=\left|I_{A}-G_{B}(I_{A})\right|$ (4)
+$\textit{L}_{identity-A}=\left|I_{A}-G_{B}(I_{A})\right|_{2}$ (4)
 and is controlled by a weight $\lambda_{i}$.
@@
 After the generator is updated, the discriminator is updated with the adversarial loss (generators frozen):
-$\textit{L}_{adv.-D_{A}}=\frac{1}{2}\left|D_{A}(I_{B})-1\right|^{2}+\frac{1}{2}\left|D_{A}(G_{A}(I_{A}))\right|^{2}$ (5)
+$\textit{L}_{adv.-D_{A}}=\frac{1}{2}\left|D_{A}(I_{B})-1\right|^{2}_{2}+\frac{1}{2}\left|D_{A}(G_{A}(I_{A}))\right|^{2}_{2}$ (5)
 For the opposite image batch $I_{B}$, the generators (B+A) and discriminator (B) are updated in the same way, as illustrated in Fig.
@@
 III-C Topology-aware Segmentation Network
-The downstream segmentation network is trained with a Perceptual Topological Loss $\mathcal{L}_{tc}$ and a Persistent Topological Loss $\mathcal{L}_{ts}$, introduced below accordingly.
+The downstream segmentation network is learned with a novel combination of Perceptual Topological Loss $\mathcal{L}_{tc}$ and a Persistent Topological Loss $\mathcal{L}_{ts}$, introduced below accordingly.
 III-C1 Perceptual Topological Loss
@@
 4, image embeddings $\mathcal{E}_{l}(\cdot)$ are computed from the predicted vessel probability map $I_{pred}$ and the ground truth $I_{GT}$, where $l$ denotes different layers in the VGG-16 network.
-In practice, we take the output of the $relu2\_2$, $relu3\_3$ and $relu4\_3$ layers.
+In practice, we take the output of the relu2_2relu2\_2, relu3_3relu3\_3 and relu4_3relu4\_3 layers.
 Then, the perceptual topological loss function $\mathcal{L}_{tc}$ is defined as:
-$\mathcal{L}_{tc}=\sum_{l}^{N_{l}}\left|\mathcal{E}_{l}(I_{pred})-\mathcal{E}_{l}(I_{GT})\right|^{2}$ (6)
+$\mathcal{L}_{tc}=\sum_{l}^{N_{l}}\left|\mathcal{E}_{l}(I_{pred})-\mathcal{E}_{l}(I_{GT})\right|^{2}_{2}$ (6)
 Figure 4: Illustration for the perceptual topological loss function.
@@
 Under each image, the digits represent the filtration value (threshold value), the number of 0-homology features and the number of 1-homology features; (c) the 0-persistent diagram; (d) the 1-persistent diagram.
-The persistent homology of an image can be described by a persistent diagram, where each point on the diagram represent one persistent feature, whose coordinate is (1--Creation Value, 1--Destruction Value).
+The persistent homology of an image can be described by a persistent diagram, where each point on the diagram represents one persistent feature, whose coordinate is (1-Creation Value, 1-Destruction Value).
 We propose to use the Spatial-Aware Topological Loss Function (SATLoss) [72] as the persistent topological loss function.
@@
 $W_{q}(\mathcal{D}(I_{pred}),\mathcal{D}(I_{GT}))=$ (7)
-$[\inf_{\eta:\mathcal{D}(I_{pred})\rightarrow\mathcal{D}(I_{GT})}\sum_{p\in\mathcal{D}(I_{pred})}s_{p}\left\|p-\eta(p))\right\|^{q}]^{\frac{1}{q}}$
+$[\inf_{\eta:\mathcal{D}(I_{pred})\rightarrow\mathcal{D}(I_{GT})}\sum_{p\in\mathcal{D}(I_{pred})}s_{p}\left\|p-\eta(p))\right\|^{q}_{2}]^{\frac{1}{q}}$
 where $s_{p}=\left\|c_{b}(p)-c_{b}(\eta(p))\right\|^{q}$ is a spatial weight that is computed between the creation cells ($c_{b}(\cdot)$) of the persistent features.
@@
 In addition, as shown in Fig.
-6, we adopt a ‘predict-refine’ strategy which first make a coarse prediction of the vessels and feed the initial prediction and the input image to the network again to refine it.
+6, we adopt a ‘predict-refine’ strategy which first makes a coarse prediction of the vessels and feeds the initial prediction and the input image to the network again to refine it.
 This method has been shown useful in both topology-aware image segmentation [62] and multiple works in deep-learning-based retinal vessel segmentation [41, 42, 43].
@@
 IV Experiments
-TABLE I: Comparison between our UVSM with SOTA finetuned methods. A green box indicates the best result and a blue box indicates the second-best result. Heidel refers to Heidelberg Spectralis.
-Dataset Modality Camera #img Method Acc. Dice Sp Se Pr F1 MCC clDice $\beta_{0}\downarrow$ $\beta_{1}\downarrow$
-$STARE$ $CF$ Topcon 20 W-Net 97.61 81.52 98.80 81.59 82.36 81.52 80.51 85.56 75.6 13.1
+TABLE I: Details of evaluation datasets
+Dataset modality camera No. images
+STARE CF Topcon 20
+HRF CF Canon 35
+ChasedDB1 CF Zeiss 28
+IOSTAR MC Heidelberg 30
+JRCFA FA Heidelberg & Optos 40
+JRCFAF FAF Heidelberg & Optos 40
+JRCIR IR Heidelberg 40
+OCTA500-3M OCTA Optovue 40
+OCTA500-6M OCTA Optovue 100
+PRIME-FP20 MC(UWF) Optos 15
+Figure 8: Qualitative comparison between our UVSM and finetuned SOTA models on multi-modality retinal images (better zoom in to view).
+Red, blue orange and green box include the original image, the ground truth label, finetuned FA-Net segmentation and the segmentation from our UVSM.
+From left to right, up to down: Topcon CF, Canon CF, Zeiss CF, Heidelberg MC, Optos FA, Heidelberg FA, Optos FAF, Heidelberg FAF, Heidelberg IR.
+Optovue OCTA 3mm, Optovue OCTA 6mm, Bottom right shows the qualitative results for Optos MC and Canon FA (no ground truth available to compare with and to train the baseline finetuned models).
+TABLE II: Comparison between our UVSM with SOTA finetuned (unsupervised domain adaptation and supervised) methods on Non - Color Fundus modality (excluding OCTA).
+The green boxes indicate the best result and the blue boxes indicate the second-best result.
+[…]
+ IOSTARIOSTAR (MC)  DomainDomain  AdaptationAdaptation Pse. Lab. 96.27 74.53 98.42 71.10 79.35 74.53 72.92 80.55 75.3 9.1
+FRR-TSNet 96.25 73.37 98.72 67.39 81.93 73.37 72.11 79.11 73.1 10.1
+MLAL 95.33 72.85 96.57 80.85 67.14 72.85 70.99 77.61 180.1 9.8
+MLAL-PDSF 95.73 74.58 96.98 81.12 69.37 74.58 72.65 80.04 138.3 6.3
+SupervisedSupervised W-Net 97.09 80.66 98.70 78.54 83.50 80.66 79.30 87.46 43.2 9.4
+WS-DMF 95.89 71.49 98.29 67.53 77.54 71.49 69.88 76.92 140.9 22.0
+GT-DLA-dsHff 96.91 79.63 98.52 78.27 81.67 79.63 78.16 86.99 43.4 7.8
+U-Net 97.01 80.50 98.52 79.72 81.80 80.50 79.04 87.08 57.4 8.7
+CS2-Net 97.05 80.59 98.59 79.27 82.49 80.59 79.17 87.06 67.2 10.3
+FA-Net 97.13 81.38 98.52 81.30 82.23 81.38 80.06 88.08 38.3 10.9
+Our UVSM 96.70 78.38 98.36 77.53 79.72 78.38 76.73 85.08 69.1 6.3
+ JRCFAJRCFA (FA)  DomainDomain  AdaptationAdaptation Pse. Lab. 88.10 6.05 94.15 5.93 7.61 6.05 0.36 3.55 251.7 34.6
+FRR-TSNet 92.71 0.60 99.56 0.33 6.01 6.04 -0.29 0.38 53.0 30.4
+MLAL 85.25 5.42 91.07 6.25 5.45 5.42 -2.07 3.02 450.1 32.2
+MLAL-PDSF 86.47 4.86 92.49 5.25 5.43 4.86 -1.88 2.73 411.6 23.5
+SupervisedSupervised W-Net 97.83 83.49 99.03 81.73 86.11 83.49 82.57 84.68 162.0 13.7
+WS-DMF 97.20 78.20 98.69 76.90 81.84 78.20 77.35 78.60 209.2 14.8
+GT-DLA-dsHff 96.38 73.14 98.17 72.25 74.44 73.14 71.31 72.16 279.1 13.2
+U-Net 97.89 83.95 99.14 81.65 87.02 83.95 83.04 84.86 160.5 13.9
+CS2-Net 97.98 84.88 99.00 84.43 85.79 84.88 83.93 86.07 157.6 11.2
+FA-Net 97.94 84.47 99.03 83.51 85.94 84.47 83.51 85.72 126.3 10.1
+Our UVSM 96.47 75.00 97.91 78.40 72.89 74.89 73.43 76.37 37.7 15.2
+ JRCFAFJRCFAF (FAF)  DomainDomain  AdaptationAdaptation Pse. Lab. 94.26 63.03 94.76 84.39 52.93 63.03 63.34 66.55 438.9 30.3
+FRR-TSNet 97.52 74.34 98.73 73.91 76.60 74.34 73.53 76.86 136.5 7.6
+MLAL 95.21 64.64 95.69 85.59 53.70 64.64 64.99 64.45 482.7 10.9
+MLAL-PDSF 95.66 66.66 96.21 85.09 56.58 66.66 66.75 67.13 425.4 19.8
+SupervisedSupervised W-Net 97.93 77.78 99.21 74.01 82.59 77.77 76.98 81.36 48.0 8.1
+WS-DMF 97.48 72.19 99.05 67.88 78.94 72.19 71.53 75.63 109.1 10.6
+GT-DLA-dsHff 97.24 70.71 98.80 67.71 74.57 70.71 69.50 74.61 62.8 7.5
+U-Net 97.89 77.66 99.13 74.94 81.29 77.66 76.81 81.03 68.6 7.2
+CS2-Net 97.93 78.11 99.14 75.34 81.64 78.12 77.24 81.46 63.4 6.9
+FA-Net 97.95 78.59 99.09 76.69 81.17 78.69 77.70 82.55 42.4 7.3
+Our UVSM 97.30 75.49 97.96 85.35 68.19 75.49 74.78 81.89 21.6 4.5
+ JRCIRJRCIR (IR)  DomainDomain  AdaptationAdaptation Pse. Lab. 95.91 71.89 97.84 71.85 73.17 71.89 70.05 71.15 86.9 8.3
+FRR-TSNet 96.06 71.95 98.21 69.38 76.13 71.95 70.30 71.02 112.1 15.3
+MLAL 94.84 69.21 96.26 76.85 64.93 69.21 67.44 68.55 239.4 14.1
+MLAL-PDSF 95.31 70.07 97.00 74.25 68.91 70.07 68.43 69.12 205.6 11.2
+SupervisedSupervised W-Net 97.07 79.79 98.85 75.25 83.35 78.80 77.52 79.19 55.3 7.0
+WS-DMF 96.32 74.49 98.10 74.37 75.87 74.49 72.88 75.67 107.9 9.8
+GT-DLA-dsHff 96.30 73.68 98.30 71.36 76.64 73.68 71.88 74.46 88.5 5.9
+U-Net 97.14 79.92 98.66 78.74 81.84 79.92 78.60 80.93 79.4 5.6
+CS2-Net 97.22 80.28 98.77 78.30 83.01 80.28 79.01 80.79 86.9 5.6
+FA-Net 97.24 80.48 98.75 78.79 82.84 80.48 79.19 81.30 49.6 5.0
+Our UVSM 96.10 74.66 97.84 75.31 72.85 73.66 71.81 76.30 28.8 5.5
+TABLE III: Comparison between our UVSM with SOTA finetuned (unsupervised domain adaptation and supervised) methods on Color Fundus modality.
+The green boxes indicate the best result and the blue boxes indicate the second-best result.
+[…]
+ STARESTARE  DomainDomain  AdaptationAdaptation Pse. Lab. 97.17 76.06 98.99 70.97 84.07 76.06 75.36 80.71 71.7 19.0
+FRR-TSNet 97.15 74.41 99.10 68.72 85.88 74.41 74.44 79.31 64.1 19.4
+MLAL 95.85 69.96 97.17 76.01 68.06 69.96 68.95 73.30 122.1 15.3
+MLAL-PDSF 96.39 73.12 97.76 76.73 71.89 73.12 71.88 76.09 129.2 15.6
+SupervisedSupervised W-Net 97.61 81.52 98.80 81.59 82.36 81.52 80.51 85.56 75.6 13.1
 WS-DMF 96.01 71.77 97.36 77.28 68.55 71.77 70.30 78.94 193.3 17.3
@@
 Our UVSM 97.48 80.51 98.69 80.77 80.78 80.52 79.31 85.21 35.1 12.6
-$ChaseDB1$ $CF$ Zeiss 28 W-Net 97.08 78.87 98.75 76.35 82.32 78.87 77.58 78.46 54.09 9.0
+ HRFHRF  DomainDomain  AdaptationAdaptation Pse. Lab. 97.00 67.36 98.83 62.73 74.10 67.36 66.38 60.19 69.6 71.5
+FRR-TSNet 97.07 66.27 99.12 58.55 78.75 66.27 66.03 59.22 75.1 74.1
+MLAL 95.65 61.96 96.94 71.07 55.87 61.96 60.52 60.77 103.8 59.0
+MLAL-PDSF 95.93 63.37 97.26 70.95 57.89 63.37 61.81 61.71 95.0 60.3
+SupervisedSupervised W-Net 96.60 64.39 98.48 61.70 68.04 64.39 62.88 60.09 260.4 59.8
+WS-DMF 94.79 58.02 96.02 71.78 49.75 58.03 56.84 56.54 945.1 51.7
+GT-DLA-dsHff 95.49 53.17 97.86 51.16 55.84 53.17 50.99 45.92 421.0 60.0
+U-Net 96.58 64.22 98.42 61.82 67.40 64.22 62.64 59.63 297.7 53.1
+CS2-Net 96.68 65.17 98.51 62.37 68.82 65.18 63.67 60.82 334.7 51.1
+FA-Net 96.67 65.80 98.40 64.36 67.87 65.80 64.23 61.97 257.7 47.5
+Our UVSM 96.60 68.50 97.78 74.70 63.87 68.50 67.15 68.26 142.8 55.5
+ $ChaseDB1$  DomainDomain  AdaptationAdaptation Pse. Lab. 96.07 71.24 98.25 68.35 75.18 71.25 69.44 70.05 68.6 6.8
+FRR-TSNet 95.94 68.93 98.53 62.93 76.63 68.93 67.26 67.17 83.5 18.3
+MLAL 94.99 68.96 96.30 78.01 62.07 68.96 66.89 67.44 203.1 23.2
+MLAL-PDSF 95.59 71.17 97.06 76.61 66.73 71.17 69.08 69.93 146.3 18.1
+SupervisedSupervised W-Net 97.08 78.87 98.75 76.35 82.32 78.87 77.58 78.46 54.09 9.0
 WS-DMF 95.54 59.62 98.81 50.31 78.94 59.62 59.92 58.15 123.6 24.1
@@
 Our UVSM 96.55 77.34 98.05 77.81 75.38 76.34 74.63 78.34 70.1 7.5
-$HRF$ $CF$ Canon 35 W-Net 96.60 64.39 98.48 61.70 68.04 64.39 62.88 60.09 260.4 59.8
-WS-DMF 94.79 58.02 96.02 71.78 49.75 58.03 56.84 56.54 945.1 51.7
-GT-DLA-dsHff 95.49 53.17 97.86 51.16 55.84 53.17 50.99 45.92 421.0 60.0
-U-Net 96.58 64.22 98.42 61.82 67.40 64.22 62.64 59.63 297.7 53.1
-CS2-Net 96.68 65.17 98.51 62.37 68.82 65.18 63.67 60.82 334.7 51.1
-FA-Net 96.67 65.80 98.40 64.36 67.87 65.80 64.23 61.97 257.7 47.5
-Our UVSM 96.60 68.50 97.78 74.70 63.87 68.50 67.15 68.26 142.8 55.5
-$IOSTAR$ $MC$ Heidel 30 W-Net 97.09 80.66 98.70 78.54 83.50 80.66 79.30 87.46 43.2 9.4
-WS-DMF 95.89 71.49 98.29 67.53 77.54 71.49 69.88 76.92 140.9 22.0
-GT-DLA-dsHff 96.91 79.63 98.52 78.27 81.67 79.63 78.16 86.99 43.4 7.8
-U-Net 97.01 80.50 98.52 79.72 81.80 80.50 79.04 87.08 57.4 8.7
-CS2-Net 97.05 80.59 98.59 79.27 82.49 80.59 79.17 87.06 67.2 10.3
-FA-Net 97.13 81.38 98.52 81.30 82.23 81.38 80.06 88.08 38.3 10.9
-Our UVSM 96.70 78.38 98.36 77.53 79.72 78.38 76.73 85.08 69.1 6.3
-$JRCFA$ $FA$ Heidel Optos 40 W-Net 97.83 83.49 99.03 81.73 86.11 83.49 82.57 84.68 162.0 13.7
-WS-DMF 97.20 78.20 98.69 76.90 81.84 78.20 77.35 78.60 209.2 14.8
-GT-DLA-dsHff 96.38 73.14 98.17 72.25 74.44 73.14 71.31 72.16 279.1 13.2
-U-Net 97.89 83.95 99.14 81.65 87.02 83.95 83.04 84.86 160.5 13.9
-CS2-Net 97.98 84.88 99.00 84.43 85.79 84.88 83.93 86.07 157.6 11.2
-FA-Net 97.94 84.47 99.03 83.51 85.94 84.47 83.51 85.72 126.3 10.1
-Our UVSM 96.47 75.00 97.91 78.40 72.89 74.89 73.43 76.37 37.7 15.2
-$JRCFAF$ $FAF$ Heidel Optos 40 W-Net 97.93 77.78 99.21 74.01 82.59 77.77 76.98 81.36 48.0 8.1
-WS-DMF 97.48 72.19 99.05 67.88 78.94 72.19 71.53 75.63 109.1 10.6
-GT-DLA-dsHff 97.24 70.71 98.80 67.71 74.57 70.71 69.50 74.61 62.8 7.5
-U-Net 97.89 77.66 99.13 74.94 81.29 77.66 76.81 81.03 68.6 7.2
-CS2-Net 97.93 78.11 99.14 75.34 81.64 78.12 77.24 81.46 63.4 6.9
-FA-Net 97.95 78.59 99.09 76.69 81.17 78.69 77.70 82.55 42.4 7.3
-Our UVSM 97.30 75.49 97.96 85.35 68.19 75.49 74.78 81.89 21.6 4.5
-$JRCIR$ IR NIR Heidel 40 W-Net 97.07 79.79 98.85 75.25 83.35 78.80 77.52 79.19 55.3 7.0
-WS-DMF 96.32 74.49 98.10 74.37 75.87 74.49 72.88 75.67 107.9 9.8
-GT-DLA-dsHff 96.30 73.68 98.30 71.36 76.64 73.68 71.88 74.46 88.5 5.9
-U-Net 97.14 79.92 98.66 78.74 81.84 79.92 78.60 80.93 79.4 5.6
-CS2-Net 97.22 80.28 98.77 78.30 83.01 80.28 79.01 80.79 86.9 5.6
-FA-Net 97.24 80.48 98.75 78.79 82.84 80.48 79.19 81.30 49.6 5.0
-Our UVSM 96.10 74.66 97.84 75.31 72.85 73.66 71.81 76.30 28.8 5.5
-Figure 8: Qualitative comparison between our UVSM and finetuned SOTA models on multi-modality retinal images (better zoom in to view).
-Red, blue orange and green box include the original image, the ground truth label, finetuned FA-Net segmentation and the segmentation from our UVSM.
-From left to right, up to down: Topcon CF, Canon CF, Zeiss CF, Heidelberg MC, Optos FA, Heidelberg FA, Optos FAF, Heidelberg FAF, Heidelberg IR.
-Bottom right shows the qualitative results for Optos MC and Canon FA (no ground truth available to compare with and to train the baseline finetuned models).
+TABLE IV: Performance on Ultra-Wide-Field retinal images, pure zero-shot performance on OCTA, and comparison with SOTA supervised methods.
+[…]
+OCTA5003MOCTA500-3M (OCTA) UNet++ 95.98 88.64 - -
+FARGO 98.12 91.68 - -
+FRNet-base 98.95 91.29 99.50 90.56
+FRNet 99.00 91.70 99.51 91.19
+Our UVSM 96.29 74.68 97.36 81.97
+OCTA5006MOCTA500-6M (OCTA) UNet++ 95.73 85.76 - -
+FARGO 98.12 89.15 - -
+FRNet-base 98.11 88.82 99.05 88.01
+FRNet 98.14 89.03 99.01 88.77
+Our UVSM 95.21 71.68 98.20 66.53
+[…]
+PRIMEFP20PRIME-FP20 (UWF-MC) LKM-UNet 58.11 41.00 57.55 51.55
+Swin-Umamba 64.99 48.43 64.36 61.93
+EM-Net 62.30 45.31 61.84 55.40
+Serp-Mamba 69.01 52.93 68.41 65.75
+Our UVSM 65.40 49.00 65.69 74.32
+TABLE V: Comparison between UVSM and general-purpose foundational segmentation models, reporting metric: Dice score.
+CF MC FA FAF IR OCTA
+SAM2 15.45 18.19 16.12 12.05 17.08 15.80
+Grounded SAM2 17.57 18.11 15.77 12.04 17.10 16.34
+MedSAM2 14.74 16.99 9.68 6.90 6.66 14.58
+Our UVSM 76.21 78.38 75.00 75.49 74.66 72.54
 IV-A Datasets
@@
 For FA, FAF and MC, 250 images are from Heidelberg camera.
-For Optos camera, the original images are ultra-wide-filed 135-degree images, which are out of the scope of this work (will be dicussed in section V-A).
-Therefore, we crop random regions from the center area of the retina to get 30-40-degree conventional field-of-view images and apply a commonly-used circular mask (Fig.
+For Optos camera, we crop random regions from the original 135 degrees to get 30-40-degree conventional field-of-view images (so that the image size is the same as the other modalities) and apply a commonly-used circular mask (Fig.
 7(a)).
@@
 We train the segmentation network using only the 20 training images from the public DRIVE [10] CF dataset.
-In practice, we discard image #14 since the vessel pattern is unusual.
-In addition, for all the images we resize them to 768$\times$768 for training.
+In practice, we discard image #14 due to the unusual vessel pattern.
+In addition, for all the images we resize them to $\times768$ for training.
 We also resized the ground truth annotation and manually corrected the error caused by interpolation.
 In addition, according to our examination, the original DRIVE ground truth has multiple topological errors such as small redundant holes in the vessel foreground.
-In our revised ground truth, we correct these errors and allow learning more accurate vessel topological features.
+In our revised ground truth, we correct these errors and allow more efficient topological learning.
 IV-A3 Evaluation
-Seven diverse datasets are used for evaluation.
+Ten diverse datasets are used for evaluation.
+Their details can be found in Table I.
 For the CF modality, we use the public STARE [83], ChaseDB1 [84] and HRF [19] datasets.
@@
 For FA, FAF and IR/NIR, we prepared three new datasets: JRCFA, JRCFAF and JRCIR.
-Each dataset has 40 768$\times$768 images, which is more than most of the commonly-used public retinal vessel segmentation (CF) datasets.
+Each dataset has 40 $\times768$ images, which is more than most of the commonly-used public retinal vessel segmentation (CF) datasets.
 The JRCFA and JRCFAF datasets both consist 20 images from Heidelberg and 20 from Optos, respectively.
@@
 All the vessel annotations are either labeled or carefully examined and refined by the retina experts in the UCSD Jacobs Retina Center (JRC).
-More information on the 7 datasets can be found in Table I.
-In addition to the seven vessel segmentation dataset, we also evaluate our model qualitatively on the Optos MC images using images from our 150K database and on the Canon FA images on the CF-FA registration dataset [81].
+Furthermore, we use an OCTA500-3M and an OCTA600-6M datasets [86] to evaluate the performance of the UVSM on OCTA in two different field-of-views (3mm and 6mm).
+For the special UWF image (technically UWF is not a modality but only refer to the field-of-view), we use a PRIME-FP20 (MC) dataset [87].
+In addition to the ten vessel segmentation dataset, we also evaluate our model qualitatively on the Optos MC images using images from our 150K database and on the Canon FA images on the CF-FA registration dataset [88].
 We provide preliminary insights in this paper (Fig.
@@
 Adam optimizer is used with a weight decay of 1e-3 and momentum of $(0.9,0.999)$.
-We train with a batch size of 4.
+We train with a batch size of 1.
 The weights $\lambda_{tc}$ and $\lambda_{ts}$ are 0.05 and 0.0002, respectively.
-All the testing images are first padded to square and resized to 768$\times$768 to pass through the translation and the segmentation network.
+All the testing images are first padded to square and resized to $\times768$ to pass through the translation and the segmentation network.
 Then the prediction is resized back to the original image size to compute the metrics with the ground truth labels.
@@
 In addition, random image rotation, shift and Gaussian noise are applied.
-Figure 9: Ablation studies on how each proposed step in the pipeline improves the performance of the universal vessel segmentation model.
-Figure 10: Qualitative comparison for how each proposed step imporves the segmentation quality (better zoom in to view). From top to bottom: CF, MC, FA, FAF, IR modality.
 IV-C Comparison with Finetuned SOTA Methods
+IV-C1 UVSM on retinal images in conventional field-of-view
 To the best of our knowledge, this work is the first universal retinal vessel segmentation model and there is no previous method that we can fairly compare with.
-Therefore, we compared with state-of-the-art (SOTA) methods which are finetuned individually on each of the 7 datasets in a fully-supervised fashion.
+Therefore, we compare with state-of-the-art (SOTA) methods which are finetuned individually on each dataset either by full-supervision or unsupervised domain adaptation.
 Since we use three new datasets and the works in the retinal vessel segmentation literature have inconsistent dataset choice, dataset split, and inconsistent choice of evaluation metrics, we prepare a new benchmark where all the methods are evaluated in a consistent way for fair comparison.
-The baseline methods are run using respective recommended training details and in a five-fold experiment that splits the whole dataset by 5 groups.
+The baseline methods are run using the recommended training details in respective papers and in a five-fold experiment that splits the whole dataset by 5 groups.
 Each fold uses one group for testing and the other four for training.
 Eventually, all images are used for testing so as to make sure that the average score is computed on the same images with our universal model (which is directly tested on all the images).
-The baseline methods include U-Net [28], CS2-Net [30], W-Net [39], FA-Net [40], GT-DLA-dsHff [31] and WS-DMF [38].
+For the unsupervised domain adaptation methods, we use the DRIVE dataset with annotation as the source domain and the per-fold training set of each dataset as the target domain for training.
+The supervised method baseline includes U-Net [31], CS2-Net [33], W-Net [42], FA-Net [43], GT-DLA-dsHff [34] and WS-DMF [41].
+For domain adapation, the baseline includes Pse.
+Lab.
+(pseudo-label-based)[42], FRR-TSNet (teacher-student network-based) [77], MLAL (adversarial learning-based) and MLAL-PDSF (hybrid adversarial learning and teacher-student) [76].
 The training and testing of the baselines are done in original image size except for the HRF dataset.
 The images in HRF dataset are too large to be trained on our 24GB GPUs.
-Therefore, we first pad the image and ground truth to square, then downsampling them to 768$\times$768 for training and testing.
+Therefore, we first pad the image and ground truth to square, then downsample them to $\times768$ for training and inference.
 The predicted vessel likelihood is then upsampled to the original image size to compare with the original ground truth and compute the metrics.
-Since the use of evaluation metrics is inconsistent in the literature, in our work we use most of the commonly-used metrics, including the pixel-wise accuracy (Acc.), the Dice score, the specificity (Sp), the sensitivity (Se), the Precision (Pr), the F1 score and the Mathew Correlation Coefficients (MCC).
-In addition, since topological learning is an important part of our work and the topological accuracy is important in retinal vessel segmentation, we also use three topological metrics which are commonly used in the segmentation of more general tubular structures.
+Since the use of evaluation metrics is inconsistent in the literature, in our work we adopt a consistent benchmark and use most of the commonly-used metrics, including the pixel-wise accuracy (Acc.), the Dice score, the specificity (Sp), the sensitivity (Se), the Precision (Pr), the F1 score and the Mathew Correlation Coefficients (MCC).
+In addition, although the main goal of topological feature learning in this work is for domain generalization, topological accuracy is still important in retinal vessel segmentation.
+Therefore, we also use three topological metrics which are commonly used in the segmentation of more general tubular structures.
 These metrics include the clDice score [65] and the 0-/1-Betti errors ($\beta_{0},\beta_{1}$) [66].
-All of the metrics are the larger the better except for the 0-/1- Betti errors, which are the smaller the better.
-For all of them reported metrics, we compute them at the original image size and averaged over all test images.
-The results in Table I, Fig.
-8 show that our proposed model, despite allowing universal segmentation, still possesses comparable performance with SOTA finetuned methods in a general sense.
-However, for some modalities there still exist non-neglectable performance gap, especially for the FA.
-As shown by an example in the second row of Fig.
+For all of the reported metrics, we compute them at the original image size and average over all test images.
+The results in Table II, Table III and Fig.
+8 show that our proposed model, not only allowing universal segmentation, but also has comparable performance with SOTA finetuned supervised methods and outperforms the finetuned unsupervised domain adaptation methods.
+However, for some modalities there still exist non-neglectable performance gap after the supervised methods, especially for the FA.
+As shown by an example in the third row of Fig.
 8, the FA modality (particularly from the Optos camera) has very high contrast between the vessel and the background.
@@
 The reasons are two-fold: (1) The image translation introduces non-avoidable detail loss; (2) The topological loss function makes the model more conservative in predicting a pixel as vessel.
-Figure 11: Comparison of the final segmentation performance between using different methods for image translation (reporting pixel-wise accuracy).
-Figure 12: Qualitative examples of the translated images using different image translation methods. (best zoom in to view)
-IV-D Ablation Studies
-IV-D1 Pipeline
+Figure 9: Qualitative results on Ultra-Wide-Field (UWF) retinal images.
+From left to right: image, translated image, UVSM segmentation, ground truth.
+Red arrows show the outlier patterns in the peripheral causing false segmentation, which is a current limitation of the UVSM.
+Figure 10: Qualitative result on how each component in the proposed pipeline acts and further comparison with other domain adaptation methods.
+Figure 11: Ablation studies on how each proposed step in the pipeline improves the performance of the universal vessel segmentation model. ‘I2I’ refers to image-to-image translation.
+IV-C2 UVSM on OCTA and UWF
+We separately evaluate OCTA modality and Ultra-Wide-Field retinal images (UWF) since they have considerably different field of views and the image feature in the OCTA modality is significantly different from the other modalities.
+In addition, vessel segmentation of these two types of retinal images has independent literature and baselines, respectively.
+For the metrics, we follow and use the metrics in [86] for OCTA modality and the metrics in [90] for UWF imaging.
+For OCTA, we compare with SOTA supervised methods: UNet++ [91], FARGO [92], FRNet-base and FRNet [86].
+We observed that the UVSM has good performance in predicting the major vessels.
+But the prediction for the thin ending vessels is usually missing (Fig.
+8).
+Generally speaking, there are still notable performance gap between the UVSM and SOTA finetuned methods (Table IV).
+But keep in mind that we only use 19 annotated CF images to train the model, while our baseline methods are finetuned on OCTA training set with highly similar data with the testing set and are over 7 times larger.
+Overall, the zero-shot results on OCTA shows the strong generalization ability of our UVSM and the effectiveness of our proposed learning methods.
+For UWF, we compare with SOTA supervised methods: LKM-UNet [93], Swin-Umamba [94], EM-Net [95] and Serp-Mamba [90].
+Experiment shows that the UVSM has strong performance in the center area of the retina but can generate more artifacts at the boundary of the image (Fig.
+9).
+The overall performance is comparable with the state-of-the-art methods (Table IV).
+The results on UWF images again shows the the strong generalization ability of our proposed model.
+Figure 12: Comparison of the final segmentation performance between using different methods for image translation. Reporting metric: Dice score.
+Figure 13: Qualitative examples of the translated images using different image translation methods. From top to bottom: MC, FA, FAF, IR, OCTA. (best zoom in to view)
+IV-D Comparison with general-purpose models
+Although great progress has been made in general-purpose segmentation models in recent years, they struggle with specialized image and structures such retinal image and curvilinear structures.
+To further highlight the need for a universal retinal vessel segmentation model, we compare with three recent foundational vision models: SAM2 [14], Grounded SAM2 [15] and the SOTA variant for general medical image segmentation-MedSAM2 [16].
+For Grounded SAM2, we use a uniform prompt word ‘vessels.’.
+For SAM2 and MedSAM2, we provide a manual point prompt with over 100 positives points on the vessel and over 20 negative points on the background for the first image of each dataset and propagate the mask to all following images.
+Due to space constraint, we report the result per modality, i.e., average the score over all images (from one or multiple datasets) from each modality.
+The UWF image is excluded from the MC modality since the extremely high resolution of UWF images makes it infeasible to run these models on our available hardware.
+The result in Table V shows that our proposed UVSM has much better performance than these general-purpose segmentation models.
+IV-E Ablation Studies
+Due to high-volume of experimental results, for all the ablation studies, we report the result per modality the same way in section IV-D.
+IV-E1 Pipeline
 In this part, we study how the main components in the proposed pipeline improve the segmentation quality.
@@
 As shown by Fig.
-9 and Fig.
-10, the performance of the model is very poor.
-After applying the image translation (CycleGAN), the segmentation quality notably improves, especially for the non-CF modalities.
-However, we still observe some results where the vessel boundary is not clear or adjacent vessels are mixing up (Fig.
+11, the performance of the model is very poor.
+Both image translation and topological learning improves the performance of the model, while topological learning plays a more dominating role.
+Optimal performance is achieved by using both I2I and topological learning.
+An interesting observation is that, for the FA modality, the vessel color is inverted and the model tends to predict the boundary of the vessels as foreground rather than the vessel itself, resulting in a ‘sandwich’ pattern.
+None other domain adaptation method can solve this issue except for the proposed topological learning (Fig.
 10).
-After further introducing the topological learning, the model makes clean and accurate segmentation.
-IV-D2 Image Translation Method
+IV-E2 Image Translation Method
 We compare CycleGAN with two most recent Diffusion model-based methods (StyleDiffusion [61] and UNSB [56]).
@@
 In general, the final segmentation performance by StyleDiffusion is the poorest (Fig.
-11).
+12).
 Furthermore, similar with the observation in [12], where style transfer method is also used for retinal vessel segmentation, the transferred images tend to include many vessel-like artifacts (Fig.
-12) introduced by the perceptual inductive bias, which will later result in many artifacts in the segmentation.
+13) introduced by the perceptual inductive bias, which will later result in many artifacts in the segmentation.
 For UNSB, the training is performed on the same dataset for training CycleGAN.
@@
 As shown in Fig.
-11, when comparing the final segmentation performance, CycleGAN demonstrates the best performance for most datasets.
+12, when comparing the final segmentation performance, CycleGAN demonstrates the best performance for most datasets.
 It is also observed that the CycleGAN translation results are visually more alike with the target CF images where the sharpness and contrast of the images are weaker than in the other modalities (Fig.
-12).
+13).
 Furthermore, for FAF where the original images are very noisy, CycleGAN also performs better in reducing the noise and generating a smoother translated image.
-Figure 13: Qualitative results for section IV-D3, From top to bottom: CF, MC, FA, FAF, IR.
-TABLE II: Comparison between different choices of topological loss functions. Light blue-conventional pixel-wise metrics, light red-topological metrics, light green-hybrid metric.
-Dataset $\mathcal{L}_{tc}$ $\mathcal{L}_{ts}$ Acc. Dice clDice $\beta_{0}$ $\beta_{1}$
-$STARE$ 97.28 79.25 83.62 65.5 16.00
-$\checkmark$ - 97.29 78.83 83.14 114.7 16.2
-- $\checkmark$ 97.08 78.20 82.86 10.7 15.4
-$\checkmark$ $\checkmark$ 97.48 80.51 85.21 35.1 12.6
-$ChaseDB1$ 96.57 77.03 78.34 92.7 10.5
-$\checkmark$ - 96.56 76.62 77.62 151.6 30.8
-- $\checkmark$ 96.34 75.32 77.22 19.1 6.4
-$\checkmark$ $\checkmark$ 96.55 77.34 78.34 70.1 7.5
-$HRF$ 96.52 67.56 66.55 160.5 66.7
-$\checkmark$ - 96.59 67.79 66.61 168.5 57.5
-- $\checkmark$ 95.99 63.43 61.67 7.4 50.9
-$\checkmark$ $\checkmark$ 96.60 68.50 68.26 70.6 55.5
-$IOSTAR$ 96.48 77.51 84.61 87.6 8.0
+Additionally, UNSB fails on the OCTA modality while StyleDiffusion and CycleGAN show better generalization ability.
+Figure 14: Qualitative results for section IV-E3, From top to bottom: CF, MC, FA, FAF, IR, OCTA
+TABLE VI: Comparison between different choices of topological loss functions.
+While the primary goal is to improve domain generalization, we still choose to report the topological metrics here since they are more relevant to the structural properties of the vessels.
+Modality $\mathcal{L}_{tc}$ $\mathcal{L}_{ts}$ Acc. Dice clDice $\beta_{0}$ $\downarrow$ $\beta_{1}$ $\downarrow$
+CFCF - - 96.72 73.57 74.64 114.7 35.5
+$\checkmark$ - 96.75 73.43 74.31 149.8 38.5
+- $\checkmark$ 96.37 71.00 72.02 12.1 27.3
+$\checkmark$ $\checkmark$ 96.80 74.38 75.75 61.9 29.0
+MCMC - - 96.48 77.51 84.61 87.6 8.0
 $\checkmark$ - 96.56 77.49 83.56 160.7 6.3
@@
 $\checkmark$ $\checkmark$ 96.70 78.38 85.08 69.1 6.3
-$JRCFA$ 90.85 40.15 36.35 233.4 17.1
+FAFA - - 90.85 40.15 36.35 233.4 17.1
 $\checkmark$ - 94.00 58.73 60.22 325.4 98.8
@@
 $\checkmark$ $\checkmark$ 96.70 75.00 76.37 37.7 15.2
-$JRCFAF$ 95.90 64.45 69.03 203.9 8.9
+FAFFAF - - 95.90 64.45 69.03 203.9 8.9
 $\checkmark$ - 97.03 73.46 77.88 149.7 8.8
@@
 $\checkmark$ $\checkmark$ 97.30 75.49 81.89 21.6 4.5
-$JRCIR$ 95.36 69.77 71.34 144.0 6.7
+IRIR - - 95.36 69.77 71.34 144.0 6.7
 $\checkmark$ - 96.01 73.48 74.66 149.2 14.6
@@
 $\checkmark$ $\checkmark$ 96.10 74.66 76.30 28.8 5.6
-IV-D3 Topological Segmentation Loss Function
-In this part we compare how the combinations of topological segmentation loss functions impact the performance of the UVSM.
-Table II and Fig.
-13 show that, both $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$ helps with overcoming the domain gap of the non-CF modalities.
+OCTAOCTA - - 89.50 33.09 33.51 83.2 7.9
+$\checkmark$ - 94.29 62.10 64.49 102.0 6.6
+- $\checkmark$ 92.22 50.89 52.45 16.8 8.3
+$\checkmark$ $\checkmark$ 95.52 72.54 76.36 13.1 7.6
+IV-E3 Topological Segmentation Loss Function
+In this part we compare how learning different combinations of topological features impact the performance of the UVSM.
+Table VI and Fig.
+14 show that, both $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$ help with overcoming the domain gap of the non-CF modalities.
 It is observed that $\mathcal{L}_{tc}$ is unable to address the inverted vessel color in FA images and tend to instead predict the boundary of the vessel, which results in a ‘sandwich’ pattern.
 In addition, using only $\mathcal{L}_{tc}$ adds saw artifacts and discontinuities to the segmentation.
-$\mathcal{L}_{ts}$ is able to address the problem of FA modality and improves the discontinuities along the vessel, however, it is observed that if only $\mathcal{L}_{ts}$ is used, the vessel boundary will be smoothed and the morphological quality of the segmentation will be poor.
+$\mathcal{L}_{ts}$ is able to address the problem of FA modality and improves the discontinuities along the vessel, however, it is observed that if only $\mathcal{L}_{ts}$ is used, the vessel boundary will be too smoothed and the morphological quality of the segmentation will be poor.
 With a combination of $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$, the saw and smoothing artifacts counteracts, and the challenge in predicting FA vessels is solved, producing the optimal results.
+TABLE VII: Comparison of using different segmentation network structures for UVSM. Reporting metric: Dice score.
+Backbone CF MC FA FAF IR OCTA
+U-Net 74.38 78.38 75.00 75.49 74.66 72.54
+FRU-Net 75.02 77.33 73.99 76.32 76.00 62.85
+MCG&BSA-Net 75.53 76.94 76.23 76.34 74.18 65.99
+RCARU-Net 76.02 78.00 75.70 76.07 75.28 60.27
+Wave-Net 75.20 76.73 76.43 76.38 75.48 71.07
+ResDoU-Net 76.01 77.89 79.13 76.15 75.49 72.71
+IV-E4 Segmentation Network Structure
+We also tested and compared multiple most recent network structures in retinal vessel segmentation as the segmentation model backbone for our UVSM, namely, FRUNet [77], MCG&BSA-Net [76], RCARU-Net [96], Wave-Net [97] and ResDoU-Net [98].
+Table VII presents that with different backbones, universal retinal vessel segmentation is still achieved, which shows the effectiveness and flexibility of our proposed learning methods.
+In addition, with some more recent backbones, the overall performance of the UVSM is further improved.
 V Conclusions
@@
 As demonstrated and discussed in the experiment section, our current universal segmentation has several limitations to improve upon.
-Firstly, although the primary vessels are well predicted, the model tends to overprune the vessels, which leads to thin vanishing vessels appearing at the peripheral area of the retina or near the macula are missing in the segmentation.
-This is caused by both the non-neglectable detail loss of the image translation process and the effect of topological loss function.
-Secondly, our current work only addresses the vessel segmentation problem for multi-modality retinal images in a most commonly-used 30-60 degree field-of-view.
-Our work does not include the ultra-wide-field 135 degrees retina images.
-Moreover, we does not consider the commonly-used 10-15 degree OCTA image, as its imaging principle is different.
-These two types of images are considerably different with the conventional modalities and requires further research to be included into the universal vessel segmentation model.
+Firstly, although the primary vessels are well predicted, the model tends to over-prune the vessels, which leads to thin vanishing vessels appearing at the peripheral area of the retina or near the macula are missing in the segmentation.
+This is caused by both the non-neglectable detail loss of the image translation process and the effect of topological loss functions.
+The effect is prominent in the OCTA and FA modalities where finer vessels are visible.
+Furthermore, the robustness of the current UVSM to outlier patterns in the peripheral area of UWF images can still be improved.
 V-B Conclusion and Future Work
-In this paper, we propose a foundational universal vessel segmentation model for multi-modality retinal images.
-Compared with the past literature of retinal vessel segmentation where (1) at most time only Color Fundus modality is studied and; (2) separate models need to be trained for different datasets (from different cameras) to achieve good performance, our method is able to perform robust vessel segmentation for all commonly-used modalities from all commonly-used cameras using only a single model and achieved comparable performance with SOTA finetuned methods.
-Although the current model still has limitations and we are working to further improve it, we believe this work is a milestone in the community of retinal image processing and understanding and marks a new direction in the research of retinal vessel segmentation.
-Our future work includes improving the universal vessel segmentation model to allow them to work for a wider range of field of views and to better predict the fine vanishing vessels.
+In this paper, we propose the universal retinal vessel segmentation task and the very first algorithm for this task, offering a generalizable and more practical solution to the retinal vessel segmentation problem.
+Compared with the past literature of retinal vessel segmentation where (1) at most time only Color Fundus modality is studied and; (2) separate models need to be finetuned for different modalities and cameras to achieve good performance, our method is able to perform modality and camera-agnostic vessel segmentation for multi-modal retinal images and achieves comparable performance with SOTA finetuned methods.
+Although the current model still has limitations and we are working to further improve it, we consider this work as a step forward in the community of retinal image processing.
+Future works include improving the universal vessel segmentation model to get better performance on fine vessels and to improve the model’s robustness to outlier patterns on the retina.

### v4

@@
 Abstract
-We identify two major limitations in the existing studies on retinal vessel segmentation: (1) Most existing works are restricted to one modality, i.e, the Color Fundus (CF).
-However, multi-modality retinal images are used every day in the study of retina and diagnosis of retinal diseases, and the study of vessel segmentation on the other modalities is scarce; (2) Even though a few works extended their experiments to limited new modalities such as the Multi-Color Scanning Laser Ophthalmoscopy (MC), these works still require finetuning a separate model for the new modality.
-The finetuning will require extra training data, which is difficult to acquire.
-In this work, we present a novel universal vessel segmentation model (UVSM) for multi-modality retinal images.
-Not only do we perform the study on a much wider range of modalities, but we also propose a universal model to segment the vessels in all these commonly-used modalities.
-Despite being much more versatile comparing with existing methods, our universal model still demonstrates comparable performance with the state-of-the-art finetuned methods.
-To the best of our knowledge, this is the first work that achieves modality-agnostic retinal vessel segmentation and also the first work that studies retinal vessel segmentation in some novel modalities.
-[^1] Our codes and model will be available at the time of publication.
-We are also actively looking for approaches to release the three new vessel segmentation datasets.
-However, since retinal images and retinal vessels contain sensitive personal information, we can not make promise at this moment.
-Index Terms:
-retina, retinal image processing, vessel segmentation, image translation, topological learning.
+We identify two major limitations in the existing studies on retinal vessel segmentation: (1) Most existing works are restricted to one modality, i.e., the Color Fundus (CF).
+However, multi-modality retinal images are used every day in the study of the retina and diagnosis of retinal diseases, and the study of vessel segmentation on other modalities is scarce; (2) Even though a few works extended their experiments to new modalities such as the Multi-Color Scanning Laser Ophthalmoscopy (MC), these works still require fine-tuning a separate model for the new modality.
+The fine-tuning will require extra training data, which is difficult to acquire.
+In this work, we present a novel universal vessel segmentation model (URVSM) for multi-modality retinal images.
+In addition to performing the study on a much wider range of image modalities, we also propose a universal model to segment the vessels in all these commonly used modalities.
+While being much more versatile compared with existing methods, our universal model also demonstrates comparable performance to the state-of-the-art fine-tuned methods.
+To the best of our knowledge, this is the first work that achieves modality-agnostic retinal vessel segmentation and the first to study retinal vessel segmentation in several novel modalities.
+[^1] Code, model and 3 new retinal vessel segmentation datasets are available at https://github.com/JRC-VPLab/URVSM [^2] This paper is accepted by IEEE Transactions on Image Processing and will be published soon.
 I Background and Introduction
 Retinal vessel segmentation is a fundamental task in retinal image processing.
-It has multiple useful applications, including retinal image registration [1, 2], human biometric identification [3], retinal artery-vein classification [4] and tree topology analysis in digital images [5].
-Clinically, the evaluation of retinal vessels is also important since the vessel features are predictive in several systemic diseases [6, 7, 8] and vessel segmentation is the key towards the automation of the evaluation.
-However, most works in retinal vessel segmentation focus on only the Color Fundus (CF) images.
-CF uses white visible light to illuminate the retina and the reflected light is used for generating the image.
-Nevertheless, although CF was the most widely-used modality in the past, it requires dilation before imaging and is uncomfortable for the patients.
-Additionally, the pathological information provided by CF is limited.
+It has numerous important applications, including retinal image registration [1, 2], human biometric identification [3], retinal artery-vein classification [4] and tree topology analysis in digital images [5].
+Clinically, the evaluation of retinal vessels is also important since the vessel features are predictive in several systemic diseases [6, 7, 8] and vessel segmentation is the key to the automation of the evaluation.
+However, most works in retinal vessel segmentation focus only on Color Fundus (CF) images.
+CF uses visible white light to illuminate the retina and the reflected light is used for generating the image.
+Nevertheless, although CF was the most widely used modality in the past, it requires dilation before imaging and is uncomfortable for patients.
+Moreover, the pathological information provided by CF is limited.
 Therefore, in more recent years, CF is used less frequently and emerging new imaging techniques become increasingly popular.
 These modalities include (Fig.
-1): (1) Multi-Color Scanning Laser Ophthalmoscopy (MC), which uses more intense laser beam and allow exposing structures in deeper layers of the retina; (2) Florescence Angiography (FA), which uses florescence dye to illuminate the retinal vessels and are commonly used in the diagnosis of diseases related to neovascularization; (3) Fundus AutoFlorescence (FAF), which does not rely on external light source but uses natural fluorescence of the retina and provides insights into the health and metabolism of the retinal pigment epithelium (RPE) and photoreceptor cells; (4) Infrared/Near-Infrared Reflectance (IR/NIR), which uses Infrared or Near-Infrared lights to illuminate the image and allows seeing deeper features on the retina such as RPE abnormalities.
-(5) Optical Coherence Tomography Angiography (OCTA), which uses low-coherence interferometry to measure the backscattered light from retinal layers.
-OCTA has a much smaller imaging area than the other modalities and has considerably different retinal vessel morphology and topology.
-Figure 1: Overview of multi-modality retinal images and the segmentation results from our UVSM.
+1): (1) Multi-Color Scanning Laser Ophthalmoscopy (MC), which uses a more intense laser beam and allows visualization of structures in deeper layers of the retina; (2) Fluorescence Angiography (FA), which uses fluorescent dye to illuminate the retinal vessels and is commonly used in the diagnosis of diseases related to neovascularization; (3) Fundus AutoFluorescence (FAF), which does not rely on an external light source but uses natural fluorescence of the retina and provides insights into the health and metabolism of the retinal pigment epithelium (RPE) and photoreceptor cells; (4) Infrared/Near-Infrared Reflectance (IR/NIR), which uses Infrared or Near-Infrared light to illuminate the image and allows visualization of deeper features on the retina, such as RPE abnormalities; (5) Optical Coherence Tomography Angiography (OCTA), which uses low-coherence interferometry to measure the backscattered light from retinal layers.
+OCTA has a much smaller imaging area than the other modalities and has a considerably different retinal vessel morphology and topology.
+Figure 1: Overview of multi-modality retinal images and the segmentation results from our URVSM.
 In each image, the upper half is the original image and the lower half is the segmentation.
@@
 Due to the complexity of the retinal vessels and the existence of many fine details, labeling is very time-consuming.
-Existing retinal vessel segmentation datasets (especially considering high label quality) are scarce and small even for the CF modality, which usually only have 20-40 images.
+Existing retinal vessel segmentation datasets (especially those with high label quality) are scarce and small, even for the CF modality, which usually only has 20-40 images.
 In addition, to the best of our knowledge, there are no publicly available vessel segmentation datasets for FAF and IR.
@@
 In this work, we prepare three high-quality datasets for FA, FAF and IR, respectively.
-Each dataset has 40 images and the vessels are carefully labeled by retina experts.
-Nevertheless, despite the availability of the new datasets, in this work, we only use them for evaluation purpose.
-The reason is that: we believe that there can never be enough annotated images to allow training a powerful universal model.
+Each dataset has 40 images, and the vessels are carefully labeled by retinal experts.
+Nevertheless, despite the availability of the new datasets, in this work, we use them only for evaluation purposes.
+The reason is that we believe that there can never be enough annotated images to allow the training of a powerful universal model.
 Instead, we need to train the model with limited existing data and seek to overcome the challenge through innovations in the algorithm.
-In fact, our segmentation network is trained with only 19 annotated CF images from the widely-used DRIVE [10] dataset.
-There are limited previous works which are extended beyond CF modality [11, 12, 13].
-In addition to lacking comprehensive modalities (compared to our approach), one still needs to finetune a separate model on the target modality where each model only works for one image modality.
-We instead aim to build a single universal model that achieves robust segmentation for all the commonly-used modalities.
-The model could serve as an important tool in cross-modality retinal image study, which becomes increasingly important in modern research of the retina.
+In fact, our segmentation network is trained with only 19 annotated CF images from the widely used DRIVE [10] dataset.
+There are limited previous works that extend beyond the CF modality [11, 12, 13].
+In addition to lacking comprehensive modalities (compared to our approach), one still needs to fine-tune a separate model on the target modality, where each model only works for one image modality.
+We instead aim to build a single universal model that achieves robust segmentation for all the commonly used modalities.
+The model could serve as an important tool in the cross-modality retinal image research, which becomes increasingly important in modern research of the retina.
 To achieve the universal segmentation with limited labeled data, we develop our model in two major steps (Fig.
 2).
-Firstly, we propose to use image translation as a preliminary domain adaptation method and train a model that translate input images from arbitrary modalities and cameras into a uniform Topcon (The Topcon TRC-50DX series is explicitly referred to as the gold standard in color fundus photography) CF image modality.
-We accessed abundant of de-identified multi-modality retinal images from multiple widely-used cameras (Topcon, Heidelberg Spectralis, Optos) at the Jacobs Retina Center at Shiley Eye Institute, University of California, San Diego (UCSD).
-Those images allow us to train a useful image translation model in a self-supervised approach.
+First, we propose to use image translation as a preliminary domain adaptation method and train a model that translates input images from arbitrary modalities and cameras into a uniform Topcon (the Topcon TRC-50DX series is explicitly referred to as the gold standard in color fundus photography) CF image modality.
+We accessed an abundance of de-identified multi-modality retinal images from multiple widely used cameras (Topcon, Heidelberg Spectralis, Optos) at the Jacobs Retina Center at Shiley Eye Institute, University of California, San Diego (UCSD).
+These images allow us to train a useful image translation model in a self-supervised manner.
 Furthermore, we adopt data augmentation to account for the slight domain difference with the other CF cameras which our center does not have, including Canon and Zeiss.
 Secondly, we train a segmentation model on the Topcon CF domain to perform the segmentation.
-However, image translation cannot fully eliminate the domain gaps between different modalities given such difficult many-to-one translation task.
+However, image translation cannot fully eliminate the domain gaps between different modalities given such a difficult many-to-one translation task.
 To solve this problem, we propose to use a topology-aware segmentation method that learns topological features of the vessels in addition to their conventional pixel-level features.
 The topology-aware segmentation methods were initially proposed to improve the topological accuracy of the segmentation of curvilinear structures.
-We go beyond their conventional applications and propose the topological learning as a novel domain adaptation method for retinal vessel segmentation, in a way that forces the segmentation to be topologically correct so when images from different domains are inputted, the model can still output robust segmentation.
-Our universal vessel segmentation model is evaluated on 10 diverse datasets (including 3 new datasets we prepared for this work) of the 6 most commonly-used retinal image modalities (CF, FA, FAF, MC, IR/NIR, OCTA) taken from 6 most commonly-used camera systems (Topcon, Canon, Zeiss, Heidelberg Spectralis, Optos, Optovue).
-Our universal model is compared with state-of-the-art methods which are finetuned on each of the 7 datasets, respectively, and is shown to achieve comparable performance.
-Extensive ablation studies were also performed to investigate how different components in the proposed pipeline affects the performance of the universal vessel segmentation.
+We go beyond their conventional applications and propose topological learning as a novel domain adaptation method for retinal vessel segmentation, in a way that forces the segmentation to be topologically correct, so when images from different domains are input, the model can still yield robust segmentation.
+Our universal vessel segmentation model is evaluated on 10 diverse datasets (including 3 new datasets we prepared for this work) of the 6 most commonly-used retinal image modalities (CF, FA, FAF, MC, IR/NIR, OCTA) taken from 6 most commonly used camera systems (Topcon, Canon, Zeiss, Heidelberg Spectralis, Optos, Optovue).
+Our universal model is compared with state-of-the-art methods that are fine-tuned on each of the 7 datasets, respectively, and is shown to achieve comparable performance.
+Extensive ablation studies were also performed to investigate how different components in the proposed pipeline affect the performance of the universal vessel segmentation.
 The main contributions of this paper are:
   • We propose the universal retinal vessel segmentation task.
-To the best of our knowledge, this is the first work that treat retinal vessel segmentation as a modality-agnostic segmentation task covering multiple imaging modalities and camera systems, reflecting the real-world need for generalizable retinal vessel segmentation solutions.
-  • We propose the first universal (retinal) vessel segmentation model (UVSM) for multi-modality retinal images based on image translation and image topological learning.
+To the best of our knowledge, this is the first work that treats retinal vessel segmentation as a modality-agnostic segmentation task covering multiple imaging modalities and camera systems, reflecting the real-world need for generalizable retinal vessel segmentation solutions.
+  • We propose the first universal (retinal) vessel segmentation model (URVSM) for multi-modality retinal images.
   • We investigate image-to-image translation between retinal images from different modalities.
-We explore different Generative Adversarial Nets (GAN)-based and Diffusion Model-based image translation methods and evaluate their performance on retinal images.
+We explore different Generative Adversarial Network (GAN)-based and Diffusion Model-based image translation methods and evaluate their performance on retinal images.
   • We propose topological feature learning as a novel domain adaptation strategy for multi-modal retinal vessel segmentation.
-We show that by learning the topological features of the vessels, the segmentation model becomes much more robust to inputs from different domains (modalities and imaging systems).
-  • We collect and annotate three high-quality vessel segmentation datasets for the FA, FAF and IR retinal images, respectively.
-  • We provide a refined version of the widely-used DRIVE dataset with improved topological accuracy in the annotation (to be discussed in section IV-A2).
+We show that by learning a combination of established topological features of the vessels, the segmentation model becomes much more robust to inputs from different domains (modalities and imaging systems).
+  • We release three new, high-quality vessel segmentation datasets for the FA, FAF, and IR retinal images, respectively.
+  • We release refined vessel annotations of the widely used DRIVE dataset with improved topological accuracy (to be discussed in section IV-A2).
 II Related Work
@@
 Current works in retinal vessel segmentation focus on segmenting vessels in Color Fundus retinal images, where various machine learning-based or non-machine learning-based methods were proposed.
-A majority of the early methods are based on the matched filter [17, 18, 19], which convolve retinal images with predefined or learned kernels and then adaptively threshold the image to obtain the vessel segmentation map.
-Another type of methods uses tracking algorithms to mathematically estimate the growth of the retinal vessel trees from seed points [20, 21].
+A majority of the early methods are based on match filtering [17, 18, 19], which convolves retinal images with predefined or learned kernels and then adaptively thresholds the image to obtain the vessel segmentation map.
+Another type of methods use tracking algorithms to mathematically estimate the growth of the retinal vessel trees from seed points [20, 21].
 The segmentation can then be derived by extending around the estimated vessel centerlines using the estimated vessel diameters.
-Some methods also adopt morphological image processing methods such as the top-hat operation to extract the vessel or as a post-processing step [22, 23].
-Furthermore, numerous traditional machine learning-based methods were also applied in retinal vessel segmentation such as the support vector machine [24], random forest [25] and Adaboost [26] for supervised methods, and Gaussian mixture model [27], Fuzzy C-Means [28] for unsupervised methods.
+Some methods also adopt morphological image processing methods such as the top-hat operation to extract the vessels or as a post-processing step [22, 23].
+Furthermore, numerous traditional machine learning-based methods were also applied in retinal vessel segmentation such as the support vector machine [24], random forest [25] and Adaboost [26] for supervised methods, and the Gaussian mixture model [27], Fuzzy C-Means [28] for unsupervised methods.
 In more recent years, deep learning-based methods dominate the literature.
-They mainly rely on the Convolutional Neural Networks (CNN) as a backbone model due to the existence of vast fine details and the spread-out characteristic of the vessels.
-An introductory work is the Deep Retinal Image Understanding (DRIU) [29], which adopts a VGG [30] network structure as encoder and assigns specialized output layers to the embedded features to segment the vessels and the optic disc from the retina.
-Based on a widely-used U-Net [31] network structure and to seek further improvement for retinal vessel segmentation, [32, 33, 34, 35, 36] proposed to add various global reasoning modules at the bottom of the network.
+They mainly rely on the Convolutional Neural Networks (CNN) as a backbone models due to the existence of vast fine details and the spread-out characteristics of the vessels.
+An introductory work is the Deep Retinal Image Understanding (DRIU) [29], which adopts a VGG [30] network structure as an encoder and assigns specialized output layers to the embedded features to segment the vessels and the optic disc from the retina.
+Based on a widely used U-Net [31] network structure and to seek further improvement for retinal vessel segmentation, [32, 33, 34, 35, 36] proposed to add various global reasoning modules at the bottom of the network.
 [35, 36, 37, 38, 34] proposed to add feature aggregation modules in the residual connection at each level of the network.
@@
 Image-to-image translation aims to transfer a source image to a target image domain while preserving its structural information.
-However, it is difficult to prepare paired retinal images from the source and target domain for training.
-Therefore, in this work we look into the unpaired image-to-image translation methods where the projection between the two domains can be learned with unpaired images with no structural correspondences.
-A representative and most-commonly used method in this task is CycleGAN [44], which introduces a cycle consistency that projects the translated image back to the source domain to force retaining the structural information in the translated image.
-Following works further introduce geometric consistency [45], mutual information regularization [46] and contrastive unpaired translation [47] to improve the computational efficiency over the cycle consistency-based methods during the training phrase.
-[48, 49] proposed one-to-many transfer schemes which allow translating images to multiple domains.
-While the above works are based on GAN backbone and achieved good image translation performance, diffusion models [50, 51, 52] have recently gained more research attention in this topic for easier training and superior image quality.
-However, the methods [53, 54, 55] are computationally intensive and were only applied on small images.
-For example, [55] combines diffusion model with adversarial learning and formulate the training in a cycle-consistency framework, which requires training two diffusion models at the same time.
-Most recently, [56] proposed an Unpaired Neuron Schrodinger Bridge (UNSB) with a novel scheme to simulate the Schrodinger Bridge, which is a random process that interpolates between two image domains, via a diffusion process.
-It has higher computational efficiency and makes application to higher-resolution images (e.g., retinal images) possible.
+However, it is difficult to prepare paired retinal images from the source and target domains for training.
+Therefore, in this work, we look into the unpaired image-to-image translation methods where the projection between the two domains can be learned with unpaired images with no structural correspondences.
+A representative and most commonly used method in this task is CycleGAN [44], which introduces a cycle consistency that projects the translated image back to the source domain to enforce retaining the structural information in the translated image.
+Following works further introduce geometric consistency [45], mutual information regularization [46], and contrastive unpaired translation [47] to improve the computational efficiency over the cycle consistency-based methods during the training phase.
+[48, 49] proposed one-to-many transfer schemes which allow the translation of images to multiple domains.
+While the above works are based on GAN backbones and achieve good image translation performance, diffusion models [50, 51, 52] have recently gained more research attention on this topic for easier training and superior image quality.
+However, the methods [53, 54, 55] are computationally intensive and were only applied to small images.
+For example, [55] combines a diffusion model with adversarial learning and formulates the training in a cycle-consistency framework, which requires training two diffusion models at the same time.
+Most recently, [56] proposed an Unpaired Neuron Schrödinger Bridge (UNSB) with a novel scheme to simulate the Schrodinger Bridge, which is a random process that interpolates between two image domains via a diffusion process.
+It has higher computational efficiency and makes the application to higher-resolution images (e.g., retinal images) possible.
 Image style transfer [57, 58, 59, 60, 61] is another method which can be seen as a special type of unpaired image-to-image translation but with a perceptual inductive bias.
-The majority of the methods use a perceptual feature space to compute a style loss and inject the style of a reference image to the input content image, while using another self-comparison content loss to retain the content information of the input in the transferred image.
+The majority of the methods use a perceptual feature space to compute a style loss and inject the style of a reference image into the input content image, while using another self-comparison content loss to retain the content information of the input in the transferred image.
 We also explore methods in this topic since style transfer is used for non-CF retinal vessel segmentation in [12, 13].
@@
 The literature mainly considers a topological loss function that extracts topological features from the ground truth and segmentation and minimizes the difference between them.
-One type of methods extracts certain features from the segmentation which are proximity to the actual topological features of the foreground objects.
-A pioneer work [62] uses an ImageNet-pretrained VGG [63, 30] network after the segmentation network.
-The perceptual features extracted from different layers of the pretrained VGG network is shown to be approximations of the topological features.
-The topological loss is then computed between the ground truth and prediction perceptual features.
+One type of methods extract certain features from the segmentation which approximate the actual topological features of the foreground objects.
+A pioneering work [62] uses an ImageNet-pretrained VGG [63, 30] network after the segmentation network.
+The perceptual features extracted from different layers of the pretrained VGG network are shown to be approximations of the topological features.
+The topological loss is then computed between the ground truth and predicted perceptual features.
 Another topological segmentation loss function [64] was proposed exclusively for the segmentation of glands.
@@
 [65] proposed to approximate the topological features of tubular structures by a relationship between the foregrounds in the ground truth and prediction and their soft-skeletons.
-Using such a relationship as a loss function is shown to strengthen the connectivity along the tubular structure thus providing better topological accuracy.
-Another type of methods uses strict topological features of the segmentation which are based on the computational topology theory and received more research attentions in recent years.
+Using such a relationship as a loss function is shown to strengthen the connectivity along the tubular structure, thus providing better topological accuracy.
+Another type of methods use strict topological features of the segmentation which are based on the computational topology theory and have received more research attention in recent years.
 The topological features are mainly based on the persistent homology theory [66, 67, 68, 69, 40, 70, 71, 72].
-A revolutionary work that first introduces this concept in image segmenation is [66], which computes the persistent features from the ground truth and prediction, matching the features using a Wasserstein distance [73] and using the distance as the loss function.
-The following works share similar idea.
-[67] also uses the Wasserstein distance but apply the method on 3D point cloud data.
+A revolutionary work that first introduced this concept in image segmentation is [66], which computes the persistent features from the ground truth and prediction, matching the features using a Wasserstein distance [73] and using the distance as the loss function.
+The following works share a similar idea.
+[67] also uses the Wasserstein distance but applies the method on 3D point cloud data.
 [70] proposed to match the longest persistent barcodes in the prediction to the ground truth.
 [40] proposed to use a Hausdorff distance [74] to match the persistent diagrams to improve the sensitivity of the matching to the outliers.
-[72] and [71] both aim to compute more precise persistent feature matching, as in earlier methods the actual matching results are highly inaccurate due to only considering the global topological features.
-[72] and [71] instead allow the loss to act more selectively on the topological features thus improve the efficiency of the topological loss function.
+[72] and [71] both aim to compute more precise persistent feature matching, as in earlier methods, the actual matching results are highly inaccurate due to only considering the global topological features.
+[72] and [71] instead allow the loss to act more selectively on the topological features thus improves the efficiency of the topological loss function.
 [71] proposed to use an Induced Matching [75] method which constructs a common ambient space between the ground truth and prediction and maps the persistent features in between to find a better matching.
 However, the computational cost of this method is very high, especially for large images, where typical retinal images usually have high resolutions.
-Alternatively, [72] proposed to leverage the spatial locations of the persistent features referencing their critical cell locations for more faithful persistent feature matching and is computationally much more efficient than [71].
-Figure 2: Pipeline for our proposed universal vessel segmentation model, which includes the 2-stage training and the inference.
+Alternatively, [72] proposes to leverage the spatial locations of the persistent features referencing their critical cell locations for more faithful persistent feature matching, and is computationally much more efficient than [71].
+Figure 2: Pipeline for our proposed universal retinal vessel segmentation model, which includes the 2-stage training and the inference.
 III Universal Retinal Vessel Segmentation Model
@@
 We discuss the overall training and inference pipeline (Fig.
-2) in this sub-section and provide more details about the major components in the pipeline in the following sub-sections.
-Given multi-modality retinal images, we first adopt a simple yet widely-used pre-processing method that takes only the green channel from the RGB image.
+2) in this subsection and provide more details about the major components in the pipeline in the following sub-sections.
+Given multi-modality retinal images, we first adopt a simple yet widely used preprocessing method that takes only the green channel from the RGB image.
 All of the training and inference are then performed on the green-channel images.
-Following pre-processing, we train a domain adaptation model for multi-modality retinal images.
-Although several domain adaptation methods were proposed for retinal vessel segmentation, such as using the pseudo-label methods [42, 76], teacher-student network [77], adversarial learning [36, 76] and asymmetrical maximum classifier discrepancy [78], they were proposed only to address the domain gap between different datasets (from different cameras) in CF images.
-Moreover, these methods still require unsupervised finetuning on the target domain images, which deviate from our goal of universal segmentation.
+Following preprocessing, we train a domain adaptation model for multi-modality retinal images.
+Although several domain adaptation methods were proposed for retinal vessel segmentation, such as using pseudo-label methods [42, 76], teacher-student networks [77], adversarial learning [36, 76] and asymmetrical maximum classifier discrepancy [78], they were proposed only to address the domain gap between different datasets (from different cameras) in CF images.
+Moreover, these methods still require unsupervised fine-tuning on the target domain images, which deviate from our goal of universal segmentation.
 Therefore, in this work we propose to use image translation as the preliminary domain adaptation method, which allows us to train a single model for all the modalities.
-The image translation model aims to translate images from all the modalities to a uniform Topcon CF modality, and is trained between equivalent number of non-CF images (FA, FAF, MC, IR/NIR) and the Topcon CF images, i.e., during the training, we treat all the non-CF modalities as one type of image.
-More information of the training dataset is provided in section IV-A1.
-Furthermore, the translation network is trained with an identity loss (will be discussed in detail in section III-B) which allows us to use a single translation network to transfer all the images to the Topcon CF domain without the need to specify whether they are CF or non-CF.
+The image translation model aims to translate images from all the modalities to a uniform Topcon CF modality, and is trained between an equivalent number of non-CF images (FA, FAF, MC, IR/NIR) and the Topcon CF images, i.e., during the training, we treat all the non-CF modalities as a single image domain.
+More information on the training dataset is provided in section IV-A1.
+Furthermore, the translation network is trained with an identity loss (will be discussed in detail in section III-B), which allows us to use a single translation network to transfer all the images to the Topcon CF domain without the need to specify whether they are CF or non-CF.
 In addition, the translation model also translates CF images from other cameras to Topcon.
@@
 We show that by learning additional topological features of the vessels, the segmentation model becomes much more robust to inputs from different domains.
-In practice, in addition to a conventional pixel-wise loss function (BCELoss), the segmentation network is jointly trained with two topological loss functions.
+In practice, in addition to a conventional pixelwise loss function (BCELoss), the segmentation network is jointly trained with two topological loss functions.
 At inference, we can disregard the modality of the image.
-Given a retinal image from an arbitrary modality, we simply need to (1) extract the green channel; (2) apply the translation network to the green-channel image; (3) apply the topology-aware segmentation network to the translated image to obtain the final segmentation.
+Given a retinal image from an arbitrary modality, we simply need to (1) extract the green channel; (2) apply the translation network to the green channel image; (3) apply the topology-aware segmentation network to the translated image to obtain the final segmentation.
 III-B Image Translation Network
-Multiple GAN-based or Diffusion model-based image translation methods are explored for our task.
-For Diffusion-based method, we experiment with StyleDiffusion [61] and Unpaired Neuron Schrodinger Bridge (UNSB) [56].
-For GAN-based method, we experiment with CycleGAN [44].
+Multiple GAN-based and diffusion model-based image translation methods are explored for our task.
+For the diffusion-based method, we experiment with StyleDiffusion [61] and Unpaired Neuron Schrödinger Bridge (UNSB) [56].
+For the GAN-based method, we experiment with CycleGAN [44].
 We will show in section IV-E2 that CycleGAN, despite being a classic method, outperforms the other two most recent methods in most datasets.
 Therefore, we pick CycleGAN as the backbone of our image translation model.
-Consequently, the following sections discuss the implementation based on CycleGAN and we refer the readers to [61, 56] for details about the other two methods and to section IV-E2 for our implementation details.
+Consequently, the following sections discuss the implementation based on CycleGAN, and we refer the readers to [61, 56] for details about the other two methods and to section IV-E2 for our implementation details.
 Figure 3: Overview of our image translation network, which is based on CycleGAN [44].
@@
 To train the translation network, we treat all the non-CF images (FA, FAF, MC, IR/NIR) as image type A and the Topcon CF images as image type B.
-At each iteration, a pair of image A batch and image B batch is selected from the training set.
+At each iteration, one batch of image type A and one batch of image type B are selected from the training set.
 As illustrated by Fig.
-3, generator A is trained to translate all the type A images (non-CF) to a fake type B (CF) and generator B is used to translate type B to a fake type A, which is a neutral non-CF domain.
-Both generator A and B are learned with adversarial learning [79, 80], where the generator and discriminator are alternatively updated.
+3, generator A is trained to translate all the type A images (non-CF) to a fake type B (CF), and generator B is used to translate type B to a fake type A, which is a neutral non-CF domain.
+Both generators A and B are learned with adversarial learning [79, 80], where the generator and discriminator are alternately updated.
 At each iteration, given the image batch $I_{A}$, the generators (A+B) are first updated with the objectives (discriminators frozen):
 $\textit{L}_{G_{A}}=\textit{L}_{adv.-G_{A}}+\lambda_{c}\textit{L}_{cycle-A}+\lambda_{i}\textit{L}_{identity-A}$ (1)
-where the adversarial loss defined as:
+where the adversarial loss is defined as:
 $\textit{L}_{adv.-G_{A}}=\left|D_{A}(G_{A}(I_{A}))-1\right|^{2}_{2}$ (2)
@@
 III-C Topology-aware Segmentation Network
-The downstream segmentation network is learned with a novel combination of Perceptual Topological Loss $\mathcal{L}_{tc}$ and a Persistent Topological Loss $\mathcal{L}_{ts}$, introduced below accordingly.
+The downstream segmentation network is learned with a novel combination of the Perceptual Topological Loss $\mathcal{L}_{tc}$ and the Persistent Topological Loss $\mathcal{L}_{ts}$, introduced below.
 III-C1 Perceptual Topological Loss
-We first propose to use a perceptual topological loss [62] where the topological information is approximated by the perceptual features encoded by an ImageNet [63] pretrained VGG-16 [30] network.
+We first propose to use a perceptual topological loss [62] where the topological information is approximated by the perceptual features encoded by an ImageNet [63]-pretrained VGG-16 [30] network.
 As illustrated by Fig.
 4, image embeddings $\mathcal{E}_{l}(\cdot)$ are computed from the predicted vessel probability map $I_{pred}$ and the ground truth $I_{GT}$, where $l$ denotes different layers in the VGG-16 network.
-In practice, we take the output of the relu2_2relu2\_2, relu3_3relu3\_3 and relu4_3relu4\_3 layers.
+In practice, we take the output of the relu2_2relu2\_2, relu3_3relu3\_3, and relu4_3relu4\_3 layers.
 Then, the perceptual topological loss function $\mathcal{L}_{tc}$ is defined as:
 $\mathcal{L}_{tc}=\sum_{l}^{N_{l}}\left|\mathcal{E}_{l}(I_{pred})-\mathcal{E}_{l}(I_{GT})\right|^{2}_{2}$ (6)
-Figure 4: Illustration for the perceptual topological loss function.
+Figure 4: Illustration of the perceptual topological loss function.
 III-C2 Persistent Topological Loss
-In 2D digital images, the topology is described by their 0- and 1-homology features, i.e., the connected components and the loops.
-The persistent homology is a method to describe a process where homology features are created and destroyed in a sequence (called ‘filtration’) of cubical complex [81] (which encodes an image and the higher-dimensional relationships between its pixels to structured cells) and its sub-complexes determined by the values of cells.
-Although it deviates from the definition and how persistent homology is computed, a simpler way to interpret persistent homology in 2D digital images is by thresholding the image from its highest pixel value to its lowest pixel value and stop at every pixel value that appears in the image.
-Each thresholded image is a binary image where we can determine the connected components and the loops inside.
+In a 2D digital image, the topology is described by its 0- and 1-homology features, i.e., the connected components and the loops.
+The persistent homology is a method to describe a process where homology features are created and destroyed in a sequence (called ‘filtration’) of a cubical complex [81] (which encodes an image and the higher-dimensional relationships between its pixels as structured cells) and its subcomplexes determined by the values of cells.
+Although it deviates from the formal definition and how persistent homology is computed, a simpler way to interpret persistent homology in 2D digital images is by thresholding the image from its highest pixel value to its lowest pixel value and stopping at every pixel value that appears in the image.
+Each thresholded image is a binary image, where we can determine the connected components and the loops inside.
 As demonstrated by the example in Fig.
-5, when the threshold value decreases, new pixels are added to the binary image, leading to creation of new homology features or destruction of old features.
+5, when the threshold value decreases, new pixels are added to the binary image, leading to the creation of new homology features or destruction of old features.
 A homology feature that ‘lives’ through the decreasing thresholds is then called a ‘persistent feature’.
-We refer the readers to [82] for more details and strict definitions.
-Figure 5: A toy example visualizing persistent homology in images: (a) a gray-scale image; (b) the filtration visualized by thresholded binary images.
-Under each image, the digits represent the filtration value (threshold value), the number of 0-homology features and the number of 1-homology features; (c) the 0-persistent diagram; (d) the 1-persistent diagram.
-The persistent homology of an image can be described by a persistent diagram, where each point on the diagram represents one persistent feature, whose coordinate is (1-Creation Value, 1-Destruction Value).
-We propose to use the Spatial-Aware Topological Loss Function (SATLoss) [72] as the persistent topological loss function.
+We refer readers to [82] for more details and strict definitions.
+Figure 5: A toy example visualizing persistent homology in images: (a) a grayscale image; (b) the filtration visualized by thresholded binary images.
+Under each image, the digits represent the filtration value (threshold value), the number of 0-homology features, and the number of 1-homology features; (c) the 0-persistent diagram; (d) the 1-persistent diagram.
+The persistent homology of an image can be described by a persistent diagram, where each point on the diagram represents one persistent feature, whose coordinates are (1-Creation Value, 1-Destruction Value).
+We propose to use a Spatial-Aware Topological Loss Function (SATLoss) [72] as the persistent topological loss function.
 After computing persistent diagrams $\mathcal{D}(I_{pred})$ and $\mathcal{D}(I_{GT})$, the persistent features are matched using a spatially-weighted Wasserstein distance [73]:
@@
 where $s_{p}=\left\|c_{b}(p)-c_{b}(\eta(p))\right\|^{q}$ is a spatial weight that is computed between the creation cells ($c_{b}(\cdot)$) of the persistent features.
-$p$ and $\eta(p)$ are a persistent feature from $I_{pred}$ and a candidate persistent feature from $I_{GT}$ to be matched with.
+$p$ and $\eta(p)$ denote a persistent feature from $I_{pred}$ and a candidate persistent feature from $I_{GT}$ to be matched.
 $q$ is the norm and in practice we take $q=2$.
@@
 $b(\eta^{*}(p))=d(\eta^{*}(p))=\frac{1}{2}(b(p)+d(p))$ (9)
-The afore-hand matching process is based on all the candidate matchings, including matching with points in $\mathcal{D}(I_{GT})$ and matching with the diagonal.
+The aforementioned matching process is based on all the candidate matchings, including matching with points in $\mathcal{D}(I_{GT})$ and matching with the diagonal.
 Figure 6: Overview of the segmentation network.
 III-C3 Loss Formulation
-Finally, the persistent topological loss and the perceptual topological loss are used with a conventional pixel-wise BCELoss:
-$\mathcal{L}_{bce}=\frac{1}{n}\sum_{i=1}^{n}y_{i}log(x_{i})+(1-y_{i})log(1-x_{i})$ (10)
-where $x_{i}$ and $y_{i}$ are the per-pixel predicted likelihood value and the ground truth value.
+Finally, the persistent topological loss and the perceptual topological loss are used with a conventional pixelwise BCELoss:
+$\mathcal{L}_{bce}=-\frac{1}{n}\sum_{i=1}^{n}y_{i}log(x_{i})+(1-y_{i})log(1-x_{i})$ (10)
+where $x_{i}$ and $y_{i}$ are the per-pixel predicted likelihood values and the ground truth value.
 $n$ is the total number of pixels in the image.
-The topological loss functions are controlled by weights $\lambda_{tc}$, $\lambda_{ts}$:
+The topological loss functions are controlled by weights $\lambda_{tc}$ and $\lambda_{ts}$:
 $\mathcal{L}_{total}=\mathcal{L}_{bce}+\lambda_{tc}\mathcal{L}_{tc}+\lambda_{ts}\mathcal{L}_{ts}$ (11)
@@
 In addition, as shown in Fig.
-6, we adopt a ‘predict-refine’ strategy which first makes a coarse prediction of the vessels and feeds the initial prediction and the input image to the network again to refine it.
-This method has been shown useful in both topology-aware image segmentation [62] and multiple works in deep-learning-based retinal vessel segmentation [41, 42, 43].
-In our method, the prediction and refinement networks share the same weight.
+6, we adopt a ‘predict-refine’ strategy, which first makes a coarse prediction of the vessels and feeds the initial prediction and the input image to the network again to refine it.
+This method has been shown to be useful in both topology-aware image segmentation [62] and multiple works in deep learning-based retinal vessel segmentation [41, 42, 43].
+In our method, the prediction and refinement networks share the same weights.
 Additionally, at both the initial prediction stage and the refinement stage, the loss functions are computed to optimize the network.
@@
 PRIME-FP20 MC(UWF) Optos 15
-Figure 8: Qualitative comparison between our UVSM and finetuned SOTA models on multi-modality retinal images (better zoom in to view).
-Red, blue orange and green box include the original image, the ground truth label, finetuned FA-Net segmentation and the segmentation from our UVSM.
-From left to right, up to down: Topcon CF, Canon CF, Zeiss CF, Heidelberg MC, Optos FA, Heidelberg FA, Optos FAF, Heidelberg FAF, Heidelberg IR.
-Optovue OCTA 3mm, Optovue OCTA 6mm, Bottom right shows the qualitative results for Optos MC and Canon FA (no ground truth available to compare with and to train the baseline finetuned models).
-TABLE II: Comparison between our UVSM with SOTA finetuned (unsupervised domain adaptation and supervised) methods on Non - Color Fundus modality (excluding OCTA).
+TABLE II: Training (one-time) and inference (per item, image size 768 $\times$ 768) time of the URVSM: translation (Trans.) + segmentation (Seg.). Device: i9-13900K CPU + 1 RTX 4090 GPU.
+Trans. Train Seg. Train Trans. Inference Seg. Inference
+6.9h 11.0h 0.15s 0.07s
+Figure 8: Qualitative comparison between our URVSM and fine-tuned SOTA models on multi-modality retinal images (best view with zooming in).
+Red, blue, orange and green box include the original image, the ground truth label, fine-tuned FA-Net segmentation and the segmentation from our URVSM.
+From left to right, top to down: Topcon CF, Canon CF, Zeiss CF, Heidelberg MC, Optos FA, Heidelberg FA, Optos FAF, Heidelberg FAF, Heidelberg IR.
+Optovue OCTA 3mm, Optovue OCTA 6mm, Bottom right shows the qualitative results for Optos MC and Canon FA (no ground truth available to compare with and to train the baseline fine-tuned models).
+TABLE III: Comparison between our URVSM and SOTA fine-tuned (unsupervised domain adaptation and supervised) methods on Non Color Fundus modality (excluding OCTA).
 The green boxes indicate the best result and the blue boxes indicate the second-best result.
@@
 FA-Net 97.13 81.38 98.52 81.30 82.23 81.38 80.06 88.08 38.3 10.9
-Our UVSM 96.70 78.38 98.36 77.53 79.72 78.38 76.73 85.08 69.1 6.3
+Our URVSM 96.70 78.38 98.36 77.53 79.72 78.38 76.73 85.08 69.1 6.3
 JRCFAJRCFA (FA) DomainDomain AdaptationAdaptation Pse. Lab. 88.10 6.05 94.15 5.93 7.61 6.05 0.36 3.55 251.7 34.6
@@
 FA-Net 97.94 84.47 99.03 83.51 85.94 84.47 83.51 85.72 126.3 10.1
-Our UVSM 96.47 75.00 97.91 78.40 72.89 74.89 73.43 76.37 37.7 15.2
+Our URVSM 96.47 75.00 97.91 78.40 72.89 74.89 73.43 76.37 37.7 15.2
 JRCFAFJRCFAF (FAF) DomainDomain AdaptationAdaptation Pse. Lab. 94.26 63.03 94.76 84.39 52.93 63.03 63.34 66.55 438.9 30.3
@@
 FA-Net 97.95 78.59 99.09 76.69 81.17 78.69 77.70 82.55 42.4 7.3
-Our UVSM 97.30 75.49 97.96 85.35 68.19 75.49 74.78 81.89 21.6 4.5
+Our URVSM 97.30 75.49 97.96 85.35 68.19 75.49 74.78 81.89 21.6 4.5
 JRCIRJRCIR (IR) DomainDomain AdaptationAdaptation Pse. Lab. 95.91 71.89 97.84 71.85 73.17 71.89 70.05 71.15 86.9 8.3
@@
 FA-Net 97.24 80.48 98.75 78.79 82.84 80.48 79.19 81.30 49.6 5.0
-Our UVSM 96.10 74.66 97.84 75.31 72.85 73.66 71.81 76.30 28.8 5.5
-TABLE III: Comparison between our UVSM with SOTA finetuned (unsupervised domain adaptation and supervised) methods on Color Fundus modality.
+Our URVSM 96.10 74.66 97.84 75.31 72.85 73.66 71.81 76.30 28.8 5.5
+TABLE IV: Comparison between our URVSM with SOTA fine-tuned (unsupervised domain adaptation and supervised) methods on Color Fundus modality.
 The green boxes indicate the best result and the blue boxes indicate the second-best result.
@@
 FA-Net 97.85 83.57 98.84 84.51 83.17 83.57 82.58 87.63 69.1 10.1
-Our UVSM 97.48 80.51 98.69 80.77 80.78 80.52 79.31 85.21 35.1 12.6
+Our URVSM 97.48 80.51 98.69 80.77 80.78 80.52 79.31 85.21 35.1 12.6
 HRFHRF DomainDomain AdaptationAdaptation Pse. Lab. 97.00 67.36 98.83 62.73 74.10 67.36 66.38 60.19 69.6 71.5
@@
 FA-Net 96.67 65.80 98.40 64.36 67.87 65.80 64.23 61.97 257.7 47.5
-Our UVSM 96.60 68.50 97.78 74.70 63.87 68.50 67.15 68.26 142.8 55.5
+Our URVSM 96.60 68.50 97.78 74.70 63.87 68.50 67.15 68.26 142.8 55.5
 $ChaseDB1$ DomainDomain AdaptationAdaptation Pse. Lab. 96.07 71.24 98.25 68.35 75.18 71.25 69.44 70.05 68.6 6.8
@@
 FA-Net 97.18 80.21 98.52 80.39 80.72 80.21 78.89 81.29 68.4 7.1
-Our UVSM 96.55 77.34 98.05 77.81 75.38 76.34 74.63 78.34 70.1 7.5
-TABLE IV: Performance on Ultra-Wide-Field retinal images, pure zero-shot performance on OCTA, and comparison with SOTA supervised methods.
+Our URVSM 96.55 77.34 98.05 77.81 75.38 76.34 74.63 78.34 70.1 7.5
+TABLE V: Performance on Ultra-Wide-Field retinal images, pure zero-shot performance on OCTA, and comparison with SOTA supervised methods.
 […]
@@
 FRNet 99.00 91.70 99.51 91.19
-Our UVSM 96.29 74.68 97.36 81.97
+Our URVSM 96.29 74.68 97.36 81.97
 OCTA5006MOCTA500-6M (OCTA) UNet++ 95.73 85.76 - -
@@
 FRNet 98.14 89.03 99.01 88.77
-Our UVSM 95.21 71.68 98.20 66.53
+Our URVSM 95.21 71.68 98.20 66.53
 […]
@@
 Serp-Mamba 69.01 52.93 68.41 65.75
-Our UVSM 65.40 49.00 65.69 74.32
-TABLE V: Comparison between UVSM and general-purpose foundational segmentation models, reporting metric: Dice score.
+Our URVSM 65.40 49.00 65.69 74.32
+TABLE VI: Comparison between URVSM and general-purpose foundational segmentation models, metric reported: Dice score.
 CF MC FA FAF IR OCTA
@@
 MedSAM2 14.74 16.99 9.68 6.90 6.66 14.58
-Our UVSM 76.21 78.38 75.00 75.49 74.66 72.54
+Our URVSM 76.21 78.38 75.00 75.49 74.66 72.54
 IV-A Datasets
@@
 However, in our previous numerous attempts we found it very difficult and time-consuming to tune the training of the image translation network (regardless of the methods) for a large dataset.
-Therefore, in this work, we use a 4k images (2k CF, 500 FA, FAF, MC, IR/NIR, respectively) dataset from the larger 150k database.
+Therefore, in this work, we use a 4k-image (2k CF, 500 FA, FAF, MC, IR/NIR, respectively) dataset from the larger 150k database.
 Each image has a resolution of 768 $\times$ 768.
 All the CF images are from the 30-degree (field-of-view) Topcon camera and all the IR/NIR images are from the Heidelberg camera.
-For FA, FAF and MC, 250 images are from Heidelberg camera.
-For Optos camera, we crop random regions from the original 135 degrees to get 30-40-degree conventional field-of-view images (so that the image size is the same as the other modalities) and apply a commonly-used circular mask (Fig.
+For FA, FAF and MC, 250 images are from the Heidelberg camera.
+For the Optos camera, we crop random regions from the original 135 degrees to get 30-40-degree conventional field-of-view images (so that the image size is the same as the other modalities) and apply a commonly used circular mask (Fig.
 7(a)).
-In the dataset, there are 250 cropped and masked Optos image for FA, MC, FAF modalities, respectively.
+In the dataset, there are 250 cropped and masked Optos images for FA, MC, FAF modalities, respectively.
 IV-A2 Segmentation Training
@@
 In addition, for all the images we resize them to $\times768$ for training.
-We also resized the ground truth annotation and manually corrected the error caused by interpolation.
+We also resized the ground truth annotations and manually corrected the errors caused by interpolation.
 In addition, according to our examination, the original DRIVE ground truth has multiple topological errors such as small redundant holes in the vessel foreground.
@@
 IV-A3 Evaluation
-Ten diverse datasets are used for evaluation.
+Ten diverse datasets were used for evaluation.
 Their details can be found in Table I.
-For the CF modality, we use the public STARE [83], ChaseDB1 [84] and HRF [19] datasets.
+For the CF modality, we use the public STARE [83], ChaseDB1 [84], and HRF [19] datasets.
 For MC modality, we use the public IOSTAR [85] dataset.
-For FA, FAF and IR/NIR, we prepared three new datasets: JRCFA, JRCFAF and JRCIR.
-Each dataset has 40 $\times768$ images, which is more than most of the commonly-used public retinal vessel segmentation (CF) datasets.
-The JRCFA and JRCFAF datasets both consist 20 images from Heidelberg and 20 from Optos, respectively.
-The JRCIR dataset consists 40 Heidelberg (the only camera for this modality) images.
-All the vessel annotations are either labeled or carefully examined and refined by the retina experts in the UCSD Jacobs Retina Center (JRC).
-Furthermore, we use an OCTA500-3M and an OCTA600-6M datasets [86] to evaluate the performance of the UVSM on OCTA in two different field-of-views (3mm and 6mm).
-For the special UWF image (technically UWF is not a modality but only refer to the field-of-view), we use a PRIME-FP20 (MC) dataset [87].
-In addition to the ten vessel segmentation dataset, we also evaluate our model qualitatively on the Optos MC images using images from our 150K database and on the Canon FA images on the CF-FA registration dataset [88].
+For FA, FAF and IR/NIR, we prepared three new datasets: JRCFA, JRCFAF, and JRCIR.
+Each dataset has 40 $\times768$ images, which is more than most of the commonly used public retinal vessel segmentation (CF) datasets.
+The JRCFA and JRCFAF datasets each consist of 20 images from Heidelberg and 20 from Optos, respectively.
+The JRCIR dataset consists of 40 Heidelberg (the only camera for this modality) images.
+All the vessel annotations are either labeled or carefully examined and refined by the retinal experts in the UCSD Jacobs Retina Center (JRC).
+Furthermore, we use an OCTA500-3M and an OCTA600-6M dataset [86] to evaluate the performance of the URVSM on OCTA in two different fields of view (3mm and 6mm).
+For the special UWF images (technically UWF refers to field-of-view rather than a distinct modality), we use the PRIME-FP20 (MC) dataset [87].
+In addition to the ten vessel segmentation datasets, we also evaluate our model qualitatively on the Optos MC images using images from our 150K database and on the Canon FA images from the CF-FA registration dataset [88].
 We provide preliminary insights in this paper (Fig.
 1, Fig.
-8) yet full evaluation requires the vessel labels (unavailable) and is left as future work.
+8), yet full evaluation requires the vessel labels (unavailable) and is left as future work.
+Additionally, we use the Massachusetts Roads [89] (road segmentation in aerial imagery) and CREMI [90] (extracellular boundary segmentation in electronic microscopy) datasets for the extended experiments on topological domain adaptation on more general curvilinear structure segmentation.
+The test set of Massachusetts Roads and all slices in volume C in CREMI were used for evaluation.
 IV-B Implementation Details
@@
 The balance between the size of the generator and discriminator is important to stably train the CycleGAN for our task.
-We use a U-Net [31] generator with 16.66M parameters and an NLP discriminator with 2.77M parameters [^2] More details on the network structures we use for translation and segmentation will be provided in our codes..
+We use a U-Net [31] generator with 16.66M parameters and an NLP discriminator with 2.77M parameters [^3] More details on the network structures we use for translation and segmentation are provided in our code..
 For training, we use a batch size of 4 and train the model for 100 epochs on the 4000 images.
@@
 The computation of persistent homology relies on the GUDHI library [91].
+The computational cost of this work is reported in Table II.
 IV-B3 Data Augmentation
@@
 For segmentation network, in addition to the aforementioned methods, we also apply random negative film followed by a contrast or sharpness adjustment.
-This is essential to allow the universal segmentation model to work for the FA modality as the vessels appears to be brighter than the background and we found no image translation method capable of stably inverting the color of the vessels.
+This is essential to allow the universal segmentation model to work for the FA modality as the vessels appear to be brighter than the background and we found no image translation method capable of stably inverting the color of the vessels.
 In addition, random image rotation, shift and Gaussian noise are applied.
-IV-C Comparison with Finetuned SOTA Methods
-IV-C1 UVSM on retinal images in conventional field-of-view
+IV-C Comparison with Fine-tuned SOTA Methods
+IV-C1 URVSM on retinal images in conventional field-of-view
 To the best of our knowledge, this work is the first universal retinal vessel segmentation model and there is no previous method that we can fairly compare with.
-Therefore, we compare with state-of-the-art (SOTA) methods which are finetuned individually on each dataset either by full-supervision or unsupervised domain adaptation.
+Therefore, we compare with state-of-the-art (SOTA) methods which are fine-tuned individually on each dataset either by full-supervision or unsupervised domain adaptation.
 Since we use three new datasets and the works in the retinal vessel segmentation literature have inconsistent dataset choice, dataset split, and inconsistent choice of evaluation metrics, we prepare a new benchmark where all the methods are evaluated in a consistent way for fair comparison.
@@
 The supervised method baseline includes U-Net [31], CS2-Net [33], W-Net [42], FA-Net [43], GT-DLA-dsHff [34] and WS-DMF [41].
-For domain adapation, the baseline includes Pse.
+For domain adaptation, the baselines include Pse.
 Lab.
@@
 The predicted vessel likelihood is then upsampled to the original image size to compare with the original ground truth and compute the metrics.
-Since the use of evaluation metrics is inconsistent in the literature, in our work we adopt a consistent benchmark and use most of the commonly-used metrics, including the pixel-wise accuracy (Acc.), the Dice score, the specificity (Sp), the sensitivity (Se), the Precision (Pr), the F1 score and the Mathew Correlation Coefficients (MCC).
-In addition, although the main goal of topological feature learning in this work is for domain generalization, topological accuracy is still important in retinal vessel segmentation.
+Since the use of evaluation metrics is inconsistent in the literature, in our work we adopt a consistent benchmark and use most of the commonly used metrics, including the pixelwise accuracy (Acc.), the Dice score, the specificity (Sp), the sensitivity (Se), the Precision (Pr), the F1 score, and the Matthews Correlation Coefficients (MCC).
+In addition, although the main goal of topological feature learning in this work is for domain generalization, topological accuracy remains crucial for retinal vessel segmentation.
 Therefore, we also use three topological metrics which are commonly used in the segmentation of more general tubular structures.
@@
 For all of the reported metrics, we compute them at the original image size and average over all test images.
-The results in Table II, Table III and Fig.
-8 show that our proposed model, not only allowing universal segmentation, but also has comparable performance with SOTA finetuned supervised methods and outperforms the finetuned unsupervised domain adaptation methods.
-However, for some modalities there still exist non-neglectable performance gap after the supervised methods, especially for the FA.
+The results in Table III, Table IV and Fig.
+8 show that our proposed model, not only allows universal segmentation, but also has comparable performance with SOTA fine-tuned supervised methods and outperforms the fine-tuned unsupervised domain adaptation methods.
+However, for some modalities there still exists non-negligible performance gap compared with the supervised methods, especially for the FA.
 As shown by an example in the third row of Fig.
@@
 Figure 9: Qualitative results on Ultra-Wide-Field (UWF) retinal images.
-From left to right: image, translated image, UVSM segmentation, ground truth.
-Red arrows show the outlier patterns in the peripheral causing false segmentation, which is a current limitation of the UVSM.
+From left to right: image, translated image, URVSM segmentation, ground truth.
+Red arrows show the outlier patterns in the peripheral causing false segmentation, which is a current limitation of the URVSM.
 Figure 10: Qualitative result on how each component in the proposed pipeline acts and further comparison with other domain adaptation methods.
 Figure 11: Ablation studies on how each proposed step in the pipeline improves the performance of the universal vessel segmentation model. ‘I2I’ refers to image-to-image translation.
-IV-C2 UVSM on OCTA and UWF
-We separately evaluate OCTA modality and Ultra-Wide-Field retinal images (UWF) since they have considerably different field of views and the image feature in the OCTA modality is significantly different from the other modalities.
-In addition, vessel segmentation of these two types of retinal images has independent literature and baselines, respectively.
+IV-C2 URVSM on OCTA and UWF
+We separately evaluate OCTA modality and Ultra-Wide-Field retinal images (UWF) since they have considerably different fields of view and the image features in the OCTA modality are significantly different from the other modalities.
+In addition, vessel segmentation of these two types of retinal images has its independent literature and established baselines.
 For the metrics, we follow and use the metrics in [86] for OCTA modality and the metrics in [92] for UWF imaging.
 For OCTA, we compare with SOTA supervised methods: UNet++ [93], FARGO [94], FRNet-base and FRNet [86].
-We observed that the UVSM has good performance in predicting the major vessels.
-But the prediction for the thin ending vessels is usually missing (Fig.
+We observed that the URVSM has good performance in predicting the major vessels, whereas the prediction for the thin ending vessels is usually missing (Fig.
 8).
-Generally speaking, there are still notable performance gap between the UVSM and SOTA finetuned methods (Table IV).
-But keep in mind that we only use 19 annotated CF images to train the model, while our baseline methods are finetuned on OCTA training set with highly similar data with the testing set and are over 7 times larger.
-Overall, the zero-shot results on OCTA shows the strong generalization ability of our UVSM and the effectiveness of our proposed learning methods.
+In general, there are still notable performance gaps between the URVSM and SOTA fine-tuned methods (Table V).
+It should be noted that we only use 19 annotated CF images to train the model, while our baseline methods are fine-tuned on OCTA training set with highly similar data with the testing set and are over 7 times larger.
+Overall, the zero-shot results on OCTA show the strong generalization ability of our URVSM and the effectiveness of our proposed learning methods.
 For UWF, we compare with SOTA supervised methods: LKM-UNet [95], Swin-Umamba [96], EM-Net [97] and Serp-Mamba [92].
-Experiment shows that the UVSM has strong performance in the center area of the retina but can generate more artifacts at the boundary of the image (Fig.
+Experiments show that the URVSM has strong performance in the center area of the retina but can generate more artifacts at the boundary of the image (Fig.
 9).
-The overall performance is comparable with the state-of-the-art methods (Table IV).
-The results on UWF images again shows the the strong generalization ability of our proposed model.
-Figure 12: Comparison of the final segmentation performance between using different methods for image translation. Reporting metric: Dice score.
-Figure 13: Qualitative examples of the translated images using different image translation methods. From top to bottom: MC, FA, FAF, IR, OCTA. (best zoom in to view)
+The overall performance is comparable with the state-of-the-art methods (Table V).
+The results on UWF images again shows the strong generalization ability of our proposed model.
+Figure 12: Comparison of the final segmentation performance between using different methods for image translation. Metric reported: Dice score.
+Figure 13: Qualitative examples of the translated images using different image translation methods. From top to bottom: MC, FA, FAF, IR, OCTA. (best viewed by zooming in)
 IV-D Comparison with general-purpose models
-Although great progress has been made in general-purpose segmentation models in recent years, they struggle with specialized image and structures such retinal image and curvilinear structures.
-To further highlight the need for a universal retinal vessel segmentation model, we compare with three recent foundational vision models: SAM2 [14], Grounded SAM2 [15] and the SOTA variant for general medical image segmentation-MedSAM2 [16].
-For Grounded SAM2, we use a uniform prompt word ‘vessels.’.
-For SAM2 and MedSAM2, we provide a manual point prompt with over 100 positives points on the vessel and over 20 negative points on the background for the first image of each dataset and propagate the mask to all following images.
-Due to space constraint, we report the result per modality, i.e., average the score over all images (from one or multiple datasets) from each modality.
+Although great progress has been made in general-purpose segmentation models in recent years, they struggle with specialized images and structures such as retinal image and curvilinear structures.
+To further highlight the need for a universal retinal vessel segmentation model, we compare with three recent foundational vision models: SAM2 [14], Grounded SAM2 [15], and the SOTA variant for general medical image segmentation: MedSAM2 [16].
+For Grounded SAM2, we use a uniform text prompt, ‘vessels.’.
+For SAM2 and MedSAM2, we provide manual point prompts with over 100 positive points on the vessels and over 20 negative points on the background for the first image of each dataset.
+The generated mask is then propagated to all subsequent images.
+Due to space constraints, we report the results per modality, i.e., average the score over all images (from one or multiple datasets) from each modality.
 The UWF image is excluded from the MC modality since the extremely high resolution of UWF images makes it infeasible to run these models on our available hardware.
-The result in Table V shows that our proposed UVSM has much better performance than these general-purpose segmentation models.
+The results in Table VI show that our proposed URVSM achieves substantially better performance than these general-purpose segmentation models.
 IV-E Ablation Studies
-Due to high-volume of experimental results, for all the ablation studies, we report the result per modality the same way in section IV-D.
+Due to the high volume of experimental results, for all the ablation studies, we report the results per modality the same way as described in section IV-D.
 IV-E1 Pipeline
@@
 All the experiments in the ablation studies adopt the same training configurations, data augmentations and network structure, unless otherwise stated.
-We first run a baseline experiment where no image translation nor topological learning is applied.
-In another word, we simply train a segmentation model on the DRIVE training set and apply the model to the multi-modality retinal images.
+We first run a baseline experiment where neither image translation nor topological learning is applied.
+In other words, we simply train a segmentation model on the DRIVE training set and apply the model to the multi-modality retinal images.
 As shown by Fig.
 11, the performance of the model is very poor.
-Both image translation and topological learning improves the performance of the model, while topological learning plays a more dominating role.
+Both image translation and topological learning improve the performance of the model, while topological learning plays a more dominant role.
 Optimal performance is achieved by using both I2I and topological learning.
 An interesting observation is that, for the FA modality, the vessel color is inverted and the model tends to predict the boundary of the vessels as foreground rather than the vessel itself, resulting in a ‘sandwich’ pattern.
-None other domain adaptation method can solve this issue except for the proposed topological learning (Fig.
+No other domain adaptation method can resolve this issue except for the proposed topological learning (Fig.
 10).
 IV-E2 Image Translation Method
-We compare CycleGAN with two most recent Diffusion model-based methods (StyleDiffusion [61] and UNSB [56]).
-The 19 translated DRIVE images are from the respective translation method to train the downstream segmentation network.
-For StyleDiffusion, we randomly pick only one CF image as the style target and reduce the training set size to the same as that in [61] because we find that on larger training set, StyleDiffusion generates severe artifacts despite our efforts to finetune the training hyperparameters for our images.
-In addition, data augmentation cannot be implemented in StyleDiffusion for its two-diffusion-stage design.
+We compare CycleGAN with the two most recent diffusion model-based methods (StyleDiffusion [61] and UNSB [56]).
+The 19 translated DRIVE images are produced by respective translation method to train the downstream segmentation network.
+For StyleDiffusion, we randomly pick only one CF image as the style target and reduce the training set size to the same as that in [61] because we find that on a larger training set, StyleDiffusion generates severe artifacts despite our efforts to fine-tune the training hyperparameters for our images.
+In addition, data augmentation cannot be implemented in StyleDiffusion because of its two-diffusion-stage design.
 In general, the final segmentation performance by StyleDiffusion is the poorest (Fig.
 12).
-Furthermore, similar with the observation in [12], where style transfer method is also used for retinal vessel segmentation, the transferred images tend to include many vessel-like artifacts (Fig.
-13) introduced by the perceptual inductive bias, which will later result in many artifacts in the segmentation.
+Furthermore, similar to the observation in [12], where style transfer methods are also used for retinal vessel segmentation, the transferred images tend to include vessel-like artifacts (Fig.
+13) introduced by the perceptual inductive bias, which in the subsequent step will also result in many artifacts in the segmentation.
 For UNSB, the training is performed on the same dataset for training CycleGAN.
@@
 12, when comparing the final segmentation performance, CycleGAN demonstrates the best performance for most datasets.
-It is also observed that the CycleGAN translation results are visually more alike with the target CF images where the sharpness and contrast of the images are weaker than in the other modalities (Fig.
+It is also observed that the CycleGAN translation results are visually more similar to the target CF images where the sharpness and contrast of the images are weaker than in the other modalities (Fig.
 13).
@@
 Figure 14: Qualitative results for section IV-E3, From top to bottom: CF, MC, FA, FAF, IR, OCTA
-TABLE VI: Comparison between different choices of topological loss functions.
+TABLE VII: Comparison between different choices of topological loss functions.
 While the primary goal is to improve domain generalization, we still choose to report the topological metrics here since they are more relevant to the structural properties of the vessels.
@@
 IV-E3 Topological Segmentation Loss Function
-In this part we compare how learning different combinations of topological features impact the performance of the UVSM.
-Table VI and Fig.
-14 show that, both $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$ help with overcoming the domain gap of the non-CF modalities.
-It is observed that $\mathcal{L}_{tc}$ is unable to address the inverted vessel color in FA images and tend to instead predict the boundary of the vessel, which results in a ‘sandwich’ pattern.
+In this part, we compare how learning different combinations of topological features impacts the performance of the URVSM.
+Table VII and Fig.
+14 show that both $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$ help overcome the domain gap of non-CF modalities.
+It is observed that $\mathcal{L}_{tc}$ is unable to address the inverted vessel color in FA images and instead tends to predict the boundary of the vessel, which results in a ‘sandwich’ pattern.
 In addition, using only $\mathcal{L}_{tc}$ adds saw artifacts and discontinuities to the segmentation.
-$\mathcal{L}_{ts}$ is able to address the problem of FA modality and improves the discontinuities along the vessel, however, it is observed that if only $\mathcal{L}_{ts}$ is used, the vessel boundary will be too smoothed and the morphological quality of the segmentation will be poor.
-With a combination of $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$, the saw and smoothing artifacts counteracts, and the challenge in predicting FA vessels is solved, producing the optimal results.
-TABLE VII: Comparison of using different segmentation network structures for UVSM. Reporting metric: Dice score.
+$\mathcal{L}_{ts}$ is able to address the problem of FA modality and reduces the discontinuities along the vessel, however, it is observed that if only $\mathcal{L}_{ts}$ is used, the vessel boundary will be too smoothed and the morphological quality of the segmentation will be poor.
+With a combination of $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$, the saw and smoothing artifacts counteract, and the challenge in predicting FA vessels is effectively mitigated, producing optimal results.
+TABLE VIII: Comparison of using different segmentation network structures for URVSM. Metric reported: Dice score.
 Backbone CF MC FA FAF IR OCTA
@@
 IV-E4 Segmentation Network Structure
-We also tested and compared multiple most recent network structures in retinal vessel segmentation as the segmentation model backbone for our UVSM, namely, FRUNet [77], MCG&BSA-Net [76], RCARU-Net [96], Wave-Net [97] and ResDoU-Net [98].
-Table VII presents that with different backbones, universal retinal vessel segmentation is still achieved, which shows the effectiveness and flexibility of our proposed learning methods.
-In addition, with some more recent backbones, the overall performance of the UVSM is further improved.
+We also tested and compared several recent network structures in retinal vessel segmentation as the segmentation model backbone for our URVSM, namely, FRUNet [77], MCG&BSA-Net [76], RCARU-Net [98], Wave-Net [99] and ResDoU-Net [100].
+Table VIII shows that with different backbones, universal retinal vessel segmentation is still achieved, which shows the effectiveness and flexibility of our proposed learning methods.
+In addition, with some more recent backbones, the overall performance of the URVSM is further improved.
+IV-F Pilot Study on Topological Domain Adaptation
+To further demonstrate the potential of topological learning as a domain adaptation strategy for more general curvilinear structure segmentation, we conducted a preliminary experiment beyond retinal vessel segmentation.
+Specifically, we directly applied the retinal vessel segmentation model trained with the proposed topological learning (corresponding to “topo” in Fig.
+11) and compared it with a baseline model trained without topological learning (“vanilla” in Fig.
+11).
+As reported in Table IX and Fig.
+15, the segmentation model trained with the joint topological learning demonstrates stronger generalization capability than its baseline counterpart.
+Figure 15: Qualitative examples for the pilot study.
+TABLE IX: Performance of retinal vessel segmentation models with/w.o. topological learning on more general curvilinear structures
+Dataset Roads CREMI
+[…]
+w.o. Topo 91.02 12.81 12.05 86.68 7.55 7.41
+with Topo 92.44 24.20 23.19 87.06 12.02 11.89
 V Conclusions
-V-A Limitations
-As demonstrated and discussed in the experiment section, our current universal segmentation has several limitations to improve upon.
-Firstly, although the primary vessels are well predicted, the model tends to over-prune the vessels, which leads to thin vanishing vessels appearing at the peripheral area of the retina or near the macula are missing in the segmentation.
-This is caused by both the non-neglectable detail loss of the image translation process and the effect of topological loss functions.
-The effect is prominent in the OCTA and FA modalities where finer vessels are visible.
-Furthermore, the robustness of the current UVSM to outlier patterns in the peripheral area of UWF images can still be improved.
-V-B Conclusion and Future Work
-In this paper, we propose the universal retinal vessel segmentation task and the very first algorithm for this task, offering a generalizable and more practical solution to the retinal vessel segmentation problem.
-Compared with the past literature of retinal vessel segmentation where (1) at most time only Color Fundus modality is studied and; (2) separate models need to be finetuned for different modalities and cameras to achieve good performance, our method is able to perform modality and camera-agnostic vessel segmentation for multi-modal retinal images and achieves comparable performance with SOTA finetuned methods.
-Although the current model still has limitations and we are working to further improve it, we consider this work as a step forward in the community of retinal image processing.
-Future works include improving the universal vessel segmentation model to get better performance on fine vessels and to improve the model’s robustness to outlier patterns on the retina.
+V-A Conclusion
+In this paper, we propose the universal retinal vessel segmentation task and the very first method for this task, offering a generalizable and more practical solution to the retinal vessel segmentation problem.
+Compared with the previous literature on retinal vessel segmentation where (1) in most cases only Color Fundus modality is studied and (2) separate models need to be fine-tuned for different modalities and cameras to achieve good performance, our method allows modality/camera-agnostic vessel segmentation for multi-modal retinal images and achieves comparable performance with SOTA fine-tuned methods.
+V-B Limitations and Future Work
+As demonstrated and discussed in the experimental section, the current version of our URVSM has several limitations that suggest directions for future improvement.
+Firstly, there remains a trade-off between universality and performance.
+While URVSM performs well on primary vessels, it tends to over-prune finer branches, leading to missed segmentation of thin peripheral vessels or capillaries near the macula.
+This contributes to a performance gap between URVSM and modality-specific SOTA models.
+The limitation arises from both the detail loss inherent to the image translation process and the regularizing effect of topological losses.
+This issue is especially evident in modalities like FA and OCTA, where capillary-level vessels could be visible.
+As a result, the current model is not suitable for capillary-level analysis, such as FAZ mapping or OCTA perfusion quantification - which we consider as an important direction for future extensions.
+Finally, although the model generalizes well to UWF images overall, its robustness to outlier patterns in the peripheral field of UWF scans can still be improved.